### v1

# Bending parameterization of one-sided degenerated Kleinian surface groups

Abstract.
It was recently proved that quasi-Fuchsian manifolds are uniquely determined by their bending laminations. This paper concerns a similar result for certain non-quasi-Fuchsian manifolds : those obtained by degenerating one end but not the other, i.e. those appearing in boundaries of Bers slices. More precisely, we show that such hyperbolic manifolds are uniquely determined by the end structure of the degenerated end and the bending lamination of the other. The end structure consists of the parabolic locus, which is a multicurve, together with ending laminations or conformal structures on each components of its complement.
1. Introduction
This paper focuses on $3$-dimensional hyperbolic manifolds (complete Riemannian $3$-manifolds locally isometric to the $3$-dimensional hyperbolic space ${\mathbb{H}}^{3}$) that are homeomorphic to $S\times{\mathbb{R}}$, where $S=S^{+}$ is a closed oriented surface of genus $\geq 2$, and $S^{-}$ is the same surface with reversed orientation. Let $\Gamma\coloneqq\pi_{1}(M)=\pi_{1}(S)$. Such a manifold is either convex-cocompact, it is then called quasi-Fuchsian, or not, in which case it can be obtained as a limit of quasi-Fuchsian ones, by the density theorem [BB04, Ohs11, NS12].
To a $3$-dimensional hyperbolic manifold $M$ homeomorphic to $S\times{\mathbb{R}}$, one can associate the action $\rho\colon\Gamma\to\rm{PSL}(2,{\mathbb{C}})\cong\rm{Isom}^{+}({\mathbb{H}}^{3})$ of its fundamental group on its universal covering $\widetilde{M}\cong{\mathbb{H}}^{3}$ by isometries, which is a discrete and faithful representation, and $M$ is homeomorphic to $M_{\rho}\coloneqq{\mathbb{H}}^{3}/\rho(\Gamma)$. There are two central objects for studying the geometry of $M$ : its conformal boundary, or boundary at infinity, and its convex core. The action $\rho$ extends to an action by conformal transformations of the conformal boundary $\partial_{\infty}{\mathbb{H}}^{3}\cong S^{2}$. Its limit set $\Lambda(\rho)$ is a closed subset of $S^{2}$, with complement its domain of discontinuity $\Omega(\rho)$. The quotient $\partial_{\infty}\rho\coloneqq\Omega(\rho)/\rho(\Gamma)$ is a finite union of Riemann surfaces of finite type [Ahl64]. It is the conformal boundary of $M$. The convex core $C(M)=C(\rho)$ is the quotient of the convex hull of $\Lambda(\rho)$ in ${\mathbb{H}}^{3}$ by the action $\rho$. It is a convex submanifold of $M$, and is actually the smallest convex submanifold whose inclusion $C(M)\xhookrightarrow{}M$ is a homotopy equivalence.
The algebraic deformation space of hyperbolic structures on $S\times{\mathbb{R}}$, denoted ${\mathcal{AH}}(S)$, is the space of discrete and faithful representations $\rho\colon\Gamma\to\rm{PSL}(2,{\mathbb{C}})$, up to conjugation. Depending on the context, an element $\rho\in{\mathcal{AH}}(S)$ is identifed with (the conjugacy class of) its image $\rho(\Gamma)\subset\rm{PSL}(2,{\mathbb{C}})$, called a Kleinian surface group, or with the hyperbolic manifold $M_{\rho}\coloneqq{\mathbb{H}}^{3}/\rho(\Gamma)$, a hyperbolic structure on $S\times{\mathbb{R}}$. We usually restrict our attention to the orientation-preserving component ${\mathcal{AH}}_{o}(S)$ of ${\mathcal{AH}}(S)$ (see Remark 2.1). [^1] This is harmless since the other component is another copy of ${\mathcal{AH}}_{o}(S)$.
The interior of ${\mathcal{AH}}_{o}(S)$ is denoted ${\mathcal{QF}}_{o}(S)$ and consists of quasi-Fuchsian representations, characterized by the compactness of their convex core. Given $\rho\in{\mathcal{QF}}_{o}(S)$, $\Omega(\rho)$ is topologically the union of two disks and $\partial_{\infty}\rho$ is a Riemann surface structure $(\operatorname{m}^{-}(\rho),\operatorname{m}^{+}(\rho))$ on $S^{-}\sqcup S^{+}$. By the simultaneous uniformization theorem of Ahlfors-Bers [Ber60], the induced map
$\operatorname{m}^{-}\times\operatorname{m}^{+}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})$
is a biholomorphic homeomorphism. Let ${\mathcal{F}}_{o}(S)\subset{\mathcal{QF}}_{o}(S)$ be the subspace of Fuchsian structures, i.e. those $\rho$ with image in $\rm{PSL}(2,{\mathbb{R}})$, up to conjugation.
Given a non-Fuchsian quasi-Fuchsian structure $\rho\in{\mathcal{QF}}_{o}(S)-{\mathcal{F}}_{o}(S)$, its convex core $C(\rho)$ is homeomorphic to $S\times[0,1]$, so its boundary $\partial C(\rho)$ is also (topologically) a copy of $S^{-}\sqcup S^{+}$, but now lying as a pleated surface inside $M_{\rho}$, i.e. a piecewise totally geodesic subsurface [Thu97, CEM06]. Its pleating or bending is recorded by a pair of measured laminations $(\operatorname{b}^{-}(\rho),\operatorname{b}^{+}(\rho))\in{\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})$ : the pleating locus constitutes a geodesic lamination and the bending angles, or limits thereof, a transverse measure. The map
$\operatorname{b}^{-}\times\operatorname{b}^{+}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})$
is continuous [KS95, Lec08]. Bonahon and Otal explicit its image in [BO04] (in the more general case of geometrically finite structures, and later generalized to the compressible case by Lecuire [Lec06]): it is the subspace ${\mathcal{ML}}^{\perp}_{<\pi}(S^{-}\sqcup S^{+})$ of filling pairs [^2] A pair $(\mu^{-},\mu^{+})\in{\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})$ is filling if, after identifying $S^{-}$ and $S^{+}$ any essential closed curve on $S$ intersects $\mu^{-}\cup\mu^{+}$. $(\mu^{-},\mu^{+})$ without closed leaves of weight $\geq\pi$. They show that it is proper onto its image, and also that $\operatorname{b}^{-}\times\operatorname{b}^{+}$ is injective over rational laminations, i.e. weighted multicurves. Bonahon also showed that $\operatorname{b}^{-}\times\operatorname{b}^{+}$ is injective in a neighborhood of Fuchsian structures inside ${\mathcal{QF}}_{o}(S)$ [Bon05]. Keen and Series proved injectivity when $S$ is the punctured torus [KS04, Ser06]. The author and Schlenker showed in [DS24] that $\operatorname{b}^{-}\times\operatorname{b}^{+}$ is injective on ${\mathcal{QF}}_{o}(S)-{\mathcal{F}}_{o}(S)$. [^3] Note that the result of Dular-Schlenker [DS24] only applies when $S$ is closed, while Bonahon’s result [Bon05] (and also Keen-Series) also covers non-closed surfaces. It follows that the map
$\operatorname{b}^{-}\times\operatorname{b}^{+}\colon{\mathcal{QF}}_{o}(S)-{\mathcal{F}}_{o}(S)\to{\mathcal{ML}}^{\perp}_{\pi}(S^{-}\sqcup S^{+})$
is a homeomorphism.
Using the same methods as in [DS24], we first give a mix of the above two parameterizations of quasi-Fuchsian structures, which will be a key step in the proof of the main result. Let ${\mathcal{ML}}_{<\pi}(S^{+})$ be the space of measured laminations on $S^{+}$ without closed leaves of weight $\geq\pi$.
Theorem (Theorem 6.1 in the text).
Consider the map $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{ML}}_{<\pi}(S^{+})$ sending a quasi-Fuchsian structure $\rho$ to the pair $(\operatorname{m}^{-}(\rho),\operatorname{b}^{+}(\rho))$ of the conformal structure of the ideal boundary of its bottom end and the bending lamination of the top boundary of its convex core. It is a homeomorphism.
The main result concerns certain hyperbolic structures on $S\times{\mathbb{R}}$ which are not quasi-Fuchsian, namely those obtained by partially degenerating one end but not the other. In this paper, we will always degenerate the bottom end and we will call them one-sided degenerated structures on $S\times{\mathbb{R}}$. Typically, such degeneration occurs at the boundary of Bers slices, i.e. $m^{+}\in{\mathcal{T}}(S^{+})$ remains fixed while $\operatorname{m}^{-}$ exits all compact subsets of ${\mathcal{T}}(S^{-})$. One can pinch a multicurve $P$ on $S^{-}$, in which case rank $1$ cusps appear in the three-manifold. Individual components of the complement $S^{-}-P$, which are cusped surfaces, can themselves be degenerated in a way that is recorded into an ending lamination. Components of $S^{-}$ that stay non-degenerated form the moderate subsurface of $S^{-}$, while the others form the immoderate subsurface. Let ${\mathcal{AH}}_{o}^{+}(S)\subset{\mathcal{AH}}_{o}(S)$ denote the space of one-sided degenerated structures, i.e. structures $\rho\in{\mathcal{AH}}_{o}(S)$ whose top end is not degenerated.
Refer to caption$\partial_{\infty}^{+}M$$\partial_{\infty}^{-}M$$C(M)$$\cong S$$C(M)$$\partial^{+}C(M)$$\partial^{-}C(M)$cuspdegenerate end
Figure 1. Schematic representation of hyperbolic manifolds homeomorphic to $S\times{\mathbb{R}}$, with $1$-dimensional “$S$-direction”, conformal boundary in red and convex core in grey. The structures are, from left to right: Fuchsian, quasi-Fuchsian, Kleinian with a cusp and Kleinian with simply-degenerated bottom end.
By the theory of quasi-conformal deformations of Kleinian groups (Ahlfors-Bers-Kra-Marden-Maskit-Sullivan [MT98, §7.55]) together with the ending lamination theorem for Kleinian surface groups by Brock, Canary and Minsky [BCM12], a Kleinian surface structure $\rho\in{\mathcal{AH}}_{o}(S)$ is entirely determined by its end invariants $(\operatorname{E}^{-}(\rho),\operatorname{E}^{+}(\rho))$, which are combinations of the data of the parabolic locus (the simple multicurve that is pinched in $S^{-}\sqcup S^{+}$ to obtain $\partial_{\infty}\rho$), and conformal structures at infinity or ending laminations on each component of the complement (see §2.7 for the notations).
In the case of one-sided degenerated structures, this says that $\rho\in{\mathcal{AH}}_{o}^{+}(S)$ is uniquely determined by a pair $(\operatorname{E}^{-}(\rho),\operatorname{m}^{+}(\rho))$, where $\operatorname{m}^{+}(\rho)$ is the top conformal structure at infinity and $\operatorname{E}^{-}(\rho)$ is the end structure of the bottom end. A consequence of the main result of this paper is the analogous statement, with $\operatorname{m}^{+}(\rho)$ replaced by the bending lamination on the top side of the convex core $\operatorname{b}^{+}(\rho)$.
Theorem.
Let $\rho\in{\mathcal{AH}}_{o}^{+}(S)$ be a one-sided degenerated Kleinian structure on $S\times{\mathbb{R}}$, and let $M_{\rho}\coloneqq{\mathbb{H}}^{3}/\rho(\pi_{1}(S))$. Then $\rho$ is uniquely determined by the pair $(\operatorname{E}^{-}(\rho),\operatorname{b}^{+}(\rho))$, i.e. by the bending lamination $\operatorname{b}^{+}(\rho)$ of the top side of the convex core of $M_{\rho}$ and by the end structure $\operatorname{E}^{-}(\rho)$ of the bottom end of $M_{\rho}$.
More precisely, let $P\subset S^{-}$ be a simple multicurve, let $\Sigma^{-}$ be a union of connected components of $S^{-}-P$ and let $L$ be a union of arational laminations on each components of $S^{-}-(P\cup\Sigma^{-})$. Denote by ${\mathcal{AH}}_{o}^{+}(S;P\cup L)\subset{\mathcal{AH}}_{o}(S)$ the subspace of one-sided degenerated structures with bottom end cusped along $P$ and having ending laminations $L$. Let <${\mathcal{ML}}^{\perp P\cup L}_{<\pi}(S^{+})\subset{\mathcal{ML}}(S^{+})$ be the subspace of measured laminations $\mu$ on $S^{+}$ without closed leaves of weight $\geq\pi$ and such that $\mu$ and $P\cup L$ fill $S-\Sigma^{-}$, i.e. any essential closed curve on $S$ intersects either $\mu\cup P\cup L$ or $\Sigma^{-}$.
Consider the map
$\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;P\cup L)\longrightarrow{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp P\cup L}_{<\pi}(S^{+})$
sending a structure $\rho$ to the pair $(\operatorname{m}^{-}(\rho),\operatorname{b}^{+}(\rho))$ of the conformal structure at infinity of the moderate subsurface of the bottom end and the bending lamination of the top side of the convex core. Then this map is a homeomorphism.
Recently, an explicit description of the combination of ending laminations and bending laminations that can occur for general Kleinian surface groups has been given by Baba and Ohshika in [BO23].
1.1. Outline of the proof
Fix a simple multicurve $P\subset S^{-}$ and a union $L$ of arational laminations in some connected components of $S^{-}-P$. Let $e^{-}=(P,L)$ and recall that ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ denotes the space of hyperbolic structures on $S\times{\mathbb{R}}$ with non-degenerated top end, and with parabolic locus $P$ and ending lamination $L$ in the bottom end. We need to prove that the map
$\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;P\cup L)\to{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}_{<\pi}^{\perp P\cup L}(S^{+})$
is continuous, proper and injective. This implies that it is a homeomorphism by invariance of domain, since both sides are topological manifolds of the same dimension.
First, we prove continuity in Section 3, see Proposition 3.4. Then, in section 4 we prove properness. The argument follows the same outline as for similar maps, considered by Bonahon-Otal [BO04], Lecuire [Lec06], Anderson-Lecuire [AL13] and Baba-Oshika [BO23]. We first show that any sequence of representations with convergent image under $\operatorname{m}^{-}\times\operatorname{b}^{+}$ must have an algebraically convergent subsequence, using the theory of degenerescence of metrics towards actions on real trees of Morgan-Shalen [MS84, Ota96]. Then, it follows from results of Anderson-Lecuire [AL13] and Bonahon-Otal and Lecuire [BO04, Lec06] that this algebraic limit must lie in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$.
To prove injectivity, we apply an argument similar to the proof of the injectivity of $\operatorname{b}^{-}\times\operatorname{b}^{+}$ for convex-cocompact structures [DS24]. More precisely, we first fix a point $m^{-}\in{\mathcal{T}}(\Sigma^{-})$ and focus on the slice ${\mathcal{AH}}_{o}^{+}(S;{\mathcal{E}}^{-})\subset{\mathcal{AH}}_{o}^{+}(S;e^{-})$ of structures $\rho$ with end structure $\operatorname{E}^{-}(\rho)={\mathcal{E}}^{-}$ consisting of $e^{-}=P\cup L$ and $m^{-}\in{\mathcal{T}}(\Sigma^{-})$. We then show that the map
$\operatorname{b}^{+}_{\infty}\colon{\mathcal{AH}}_{o}^{+}(S;{\mathcal{E}}^{-})\to{\mathcal{ML}}_{<\pi}^{\perp e^{-}}(S^{+})$
is injective, by exhibiting it as a limit of homeomorphisms $\operatorname{b}^{+}_{t}$, $t\in[0,\infty)$. Combined with the fact that $\operatorname{b}^{+}_{\infty}$ has compact real-analytic fibres and using a theorem of Finney [Fin67], this implies that $\operatorname{b}^{+}_{\infty}$ is injective.
[…]
${{[0,\infty)\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{QF}}_{o}(S)}}$${{{\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})}}$${{[0,\infty]\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{AH}}_{o}^{+}(S)}}$${{{\mathcal{ML}}(S^{+})}}$$\scriptstyle{r\times\rm{incl}}$$\scriptstyle{\operatorname{U}}$$\scriptstyle{\operatorname{b}^{+}\times\operatorname{b}^{-}}$$\scriptstyle{\rm{proj}_{2}}$$\scriptstyle{\overline{r}\times\rm{incl}}$$\scriptstyle{\operatorname{B}^{+}}$¯¯$\scriptstyle{\overline{\operatorname{U}}}$$\scriptstyle{\operatorname{b}^{+}}$
Let $\operatorname{B}^{+}\colon[0,\infty]\times{\mathcal{T}}(S^{+})\to{\mathcal{ML}}(S^{+})$ be the composition $\operatorname{B}^{+}\coloneqq\operatorname{b}^{+}\circ\,\overline{\operatorname{U}}\circ(\overline{r}\times\rm{incl})$, and for each $t\in[0,\infty]$, write $\operatorname{b}_{t}^{+}\coloneqq\operatorname{B}^{+}(t,\cdot)\colon{\mathcal{T}}(S^{+})\to{\mathcal{ML}}(S^{+})$.
To show that $\operatorname{b}_{\infty}^{+}$ is injective, we need the following ingredients :
  1. I
    The map $\operatorname{B}^{+}$ is continuous (Corollary 3.5), hence $\operatorname{b}^{+}_{\infty}$ is the continuous limit of $\operatorname{b}^{+}_{t}$ when $t\to\infty$ ;
  2. II
    The map $\operatorname{b}^{+}_{\infty}$ is proper (§4) ;
  3. III
    The map $\operatorname{b}^{+}_{\infty}$ has real-analytic fibres (§5) ;
  4. IV
    For each $t\in[0,\infty)$, the map $\operatorname{b}^{+}_{t}$ is injective (by invariance of domain, it is then an open embedding) (§6).
Combining (I) and (IV), Theorem 6.7, which is a variant of Finney’s theorem [Fin67, DS24], implies that the compact ”topologically nice” fibres of $\operatorname{b}^{+}_{\infty}$ are contractible. Any fibre $F$ of $\operatorname{b}^{+}_{\infty}$ is compact by (II) and ”nice” by (III). All in all, $F$ is a contractible compact real-analytic space. But a compact real-analytic space $F$ has a fundamental class modulo $2$ [BH61, Sul71], i.e. a nonzero homology class in $H_{\dim F}(F;{\mathbb{Z}}/2)$. Since $F$ is contractible, we must have $\dim F=0$, which means that $F$ is a point. This shows that $\operatorname{b}^{+}_{\infty}$ is injective.
In Section 2 the necessary background is recalled. Sections 3-6 are devoted to proving the above four ingredients, with the conclusion of the proof of the main result at the end of Section 6.
1.2. Further directions
A natural question is whether any structure in ${\mathcal{AH}}_{o}(S)$ is uniquely determined by the combination of its bending and ending laminations. The main theorem of this paper implies that it is the case when the top side is non-degenerated and the bottom moderate subsurface has no quasi-conformal deformations, i.e. $\Sigma^{-}$ is empty or a union of thrice-punctured spheres. But this is still far from a complete answer.
The essential difficulty lies in proving that any geometrically finite structure on $S\times{\mathbb{R}}$ is uniquely determined by its pair of bending laminations. This is expected to hold, and it was proved in the case of rational bending laminations by Bonahon and Otal [BO04]. But it is not yet known in the general case, where our strategy fails.
If geometrically finite structures are proven to be parameterized by their bending laminations, then one can prove that the same holds for all structures in ${\mathcal{AH}}_{o}(S)$. Indeed, any structure in ${\mathcal{AH}}_{o}(S)$ can be obtained as a limit of geometrically finite structures with the same parabolic locus, and the same argument as the one of the present main result can be applied, combined with the properness result of Baba and Ohshika [BO23].
Acknowledgements
I wish to thank Tommaso Cremaschi, Cyril Lecuire and Jean-Marc Schlenker for helpful discussions. I was supported by FNR AFR grant 18890152.
2. Background
The main references used are the books of Matsuzaki-Taniguchi [MT98], Canary-Epstein-Marden [CEM06] and Marden [Mar16].
Unless stated otherwise, $S=S^{+}$ denotes a closed oriented surface of genus $\geq 2$, and $S^{-}$ is the same surface with reversed orientation. Let $\Gamma\coloneqq\pi_{1}(S)$. The space of representations of $\Gamma$ into $\rm{PSL}(2,{\mathbb{C}})\cong\rm{Isom}^{+}({\mathbb{H}}^{3})$, modulo conjugation, is denoted ${\mathcal{X}}(S)$. Let ${\mathcal{AH}}(S)\subset{\mathcal{X}}(S)$ denote the subspace of discrete and faithful representations. Any $\rho\in{\mathcal{AH}}(S)$ determines a hyperbolic manifold $M_{\rho}\coloneqq{\mathbb{H}}^{3}/\rho(\Gamma)$, which is homeomorphic to $S\times{\mathbb{R}}$ by Bonahon’s tameness theorem [Bon86]. Thus ${\mathcal{AH}}(S)$ is also called the algebraic deformation space of marked hyperbolic structures on $S\times{\mathbb{R}}$.
Remark 2.1.
Let $\rho\in{\mathcal{AH}}(S)$. Under the homeomorphism $M_{\rho}\cong S\times{\mathbb{R}}$, $M_{\rho}$ has the same orientation as $S^{+}\times{\mathbb{R}}$, or the reversed orientation. In fact, ${\mathcal{AH}}(S)$ has two connected components, characterized by this property [Mar16, §5.2.1]. The orientation-preserving component of ${\mathcal{AH}}(S)$, denoted ${\mathcal{AH}}_{o}(S)$ is the component consisting of those $\rho\in{\mathcal{AH}}(S)$ for which $M_{\rho}$ is homeomorphic to $S^{+}\times{\mathbb{R}}$ in an orientation-preserving way. When we speak of the bottom and top ends of a Kleinian surface manifold, we implicitely restrict our attention to elements of the orientation-preserving component.
Given $\rho\in{\mathcal{AH}}(S)$ and $c\subset S$ a curve on $S$, we denote by $\ell_{\rho}(c)$ the length of a geodesic representative of $c$ inside $M_{\rho}$, or zero if $c$ has no such geodesic representative. Equivalently, if $\gamma\in\pi_{1}(S)$ represents $c$, then $\ell_{\rho}(c)$ is the translation distance of the isometry $\rho(\gamma)$ on ${\mathbb{H}}^{3}$.
2.2. Laminations and trees
Let ${\mathcal{GL}}(S)$ be the space of geodesic laminations on $S$, i.e. closed subsets that are the union of disjoint simple geodesics (its leaves), endowed with the Chabauty topology (which coincides with the Hausdorff topology and geometric topology [CEM06, §I.3.1]). A geodesic lamination is said to be minimal or maximal when it is so with respect to inclusion (among non-empty laminations). A minimal geodesic lamination is either a simple closed curve or it is irrational, in which case it contains uncountably many geodesics and any half-leaf is dense in it. A leaf of a geodesic lamination $L$ is recurrent if it is contained in a minimal sublamination. A geodesic lamination is recurrent if all its leaves are. A geodesic lamination is rational if it is a simple multicurve, i.e. a finite union of simple closed curves, and arational if it is minimal and maximal (in which case it is irrational and each complement of its complement is either simply-connected or an annulus around a cusp). The space ${\mathcal{GL}}(S)$ is compact [CEM06, I.4.1.7].
A transverse measure on a geodesic lamination $L\in{\mathcal{GL}}(S)$ is a map $\mu$ associating to each closed arc transverse to $L$ a number $\mu(k)\in{\mathbb{R}}_{\geq 0}$ (also denoted $i(\mu,k)$) which is invariant under transverse isotopy of arcs, countably additive, and with full support, i.e. $\mu(k)>0$ if and only if $k\cap L\not=\emptyset$. Let ${\mathcal{ML}}(S)$ denote the space of measured lamination on $S$, endowed with the weak topology. The geodesic laminations that admit a transverse measure (with full support) are precisely the recurrent laminations [PH92, §1.7]. A minimal geodesic lamination admits a unique transverse measure up to scaling. See [PH92] for more details on measured laminations, and a description of a piecewise linear structure on ${\mathcal{ML}}(S)$ using train tracks. By Thurston earthquake theorem [Thu86], the space ${\mathcal{ML}}(S)$ is homeomorphic to ${\mathcal{T}}(S)$. See also [CEM06], [Ota96, Appendix], [MT98, §6.1].
Denote by $i\colon{\mathcal{ML}}(S)\times{\mathcal{ML}}(S)\to{\mathbb{R}}_{\geq 0}$ the intersection form between measured laminations, which continuously extends both the intersection number between weighted multicurve and the measure of transverse arcs.
To a measured lamination $\mu$ on $S$ one can associate its dual real tree ${\mathcal{T}}_{\mu}$, together with an action $\rho$ of $\pi_{1}(S)$ on ${\mathcal{T}}_{\mu}$ by isometries [Ota96, §8.2]. For $\gamma\in\pi_{1}(S)$, the translation distance of $c$ on ${\mathcal{T}}_{\mu}$ is precisely $i(\mu,c)$. A sequence of representations $(\rho_{n})_{n\in{\mathbb{N}}}\subset{\mathcal{AH}}(S)$ is said to converge to an action $\rho$ of $\pi_{1}(S)$ on a real tree ${\mathcal{T}}_{\mu}$ if there is a sequence of positive real numbers $\epsilon\searrow 0$ such that $\epsilon_{n}\ell_{\rho_{n}}(\gamma)\to i(\mu,\gamma)$ for all $\gamma\in\pi_{1}(S)$.
Morgan and Shalen proved that actions on real trees arise naturally as limits of divergent sequences of discrete and faithful representations of $\pi_{1}(S)$ in $\rm{Isom}^{+}({\mathbb{H}}^{3})$, as stated below [MS84]. The fact that real trees arising in this way are in fact dual to measured laminations is due to Skora [Sko96]. See [Ota96] for a detailed exposition.
Theorem 2.3 (Morgan-Shalen [MS84, Sko96, Ota96]).
Let $(\rho_{n})_{n\in{\mathbb{N}}}$ be a sequence in ${\mathcal{AH}}(S)$. Then, up to extracting a subsequence, either :
  1. (1)
    The sequence $(\rho_{n})_{n}$ converges in ${\mathcal{AH}}(S)$, or ;
  2. (2)
    The sequence $(\rho_{n})_{n}$ converges to an action $\rho$ of $\pi_{1}(S)$ on a real tree ${\mathcal{T}}_{\mu}$, where $\mu$ is a measured lamination on $S$.
A measured lamination $\alpha$ is said to be realized in the real tree ${\mathcal{T}}_{\mu}$ if all its leaves intersect $\mu$ transversely [Ota96, §3.1]. See [MS84], [Sko96] and [Ota96] for more details.
2.4. Pleated surfaces
Let $M$ be a hyperbolic three-manifold, $S$ a complete hyperbolic surface (not necessarily compact) and $L$ a geodesic lamination on $S$. A pleated surface with pleating locus $L$ is a continuous map $f\colon S\to M$ which is a totally geodesic immersion on $S-L$ and on $L$. In general, the flat pieces of $S-L$ are bent along the pleating locus, with an angle recorded into a bending cocycle $\beta$ which is an element of the space ${\mathcal{H}}(L;{\mathbb{R}}/2\pi{\mathbb{Z}})$ of ${\mathbb{R}}/2\pi{\mathbb{Z}}$-valued cocycles along $L$ [Bon96] (which is a torus whose dimension is determined by $S$ and $L$). When $f$ is locally convex, this bending cocycle defines a transverse measure with support contained in $L$, called the bending or pleating (measured) lamination of $f$. The induced path metric on $S$ inside $M$ is hyperbolic and defines a point $h\in{\mathcal{T}}(S)$, which can also be encoded in a shearing cocycle $\sigma\in{\mathcal{H}}(L;{\mathbb{R}})$. The boundary of the convex core $C(\rho)$ of $\rho\in{\mathcal{AH}}(S)$ is an embedded convex pleated surface [CEM06, II.1.11,12]. In particular, its bending measured lamination has weight $<\pi$ along closed leaves.
Now, let $L$ be a maximal geodesic lamination. Let $\rho\in{\mathcal{AH}}(S)$. If there exists a pleated surface inside $M_{\rho}$ with pleating locus $L$ (we say that $L$ is realized in $M_{\rho}$ or by $\rho$), then it is unique and the shear-bend cocycle provides a parameterization of such representations. More generally, one can consider abstract pleated surfaces $(\widetilde{f},\pi_{1}(S),\rho)$ where $\widetilde{f}\colon\widetilde{S}\to{\mathbb{H}}^{3}$ is a pleated surface inside ${\mathbb{H}}^{3}$ which is equivariant with respect to the (not necessarily discrete) representation $\rho\in{\mathcal{X}}(S)$. Given a maximal geodesic lamination, let ${\mathcal{R}}(L)$ denote the open subspace of ${\mathcal{X}}(S)$ consisting of those representations realizing $L$.
Theorem 2.5 (Bonahon [Bon96, Theorem D]).
Let $L$ be a maximal geodesic lamination. The map ${\mathcal{R}}(L)\to{\mathcal{H}}(L;{\mathbb{C}}/2\pi i{\mathbb{Z}})$ sending a representation $\rho$ realizing $L$ to its shear-bend cocycle $\sigma(\rho)+i\beta(\rho)$ is a biholomorphic homeomorphism onto an open subset of ${\mathcal{H}}(L;{\mathbb{C}}/2\pi i{\mathbb{Z}})$.
Using pleated surface one can also speak of the length of a realizable measured lamination in a Kleinian surface manifold. Consider $\lambda\in{\mathcal{ML}}(S)$ and $\rho\in{\mathcal{AH}}(S)$. If $\rho\in{\mathcal{R}}(|\lambda|)$, i.e. $M_{\rho}$ contains a pleated surface $P$ pleated along the support $|\lambda|$ of $\lambda$, then the length $\ell_{\rho}(\lambda)$ is defined as the length of $\lambda$ on $P$. When $\lambda$ is a closed curve, this coincides with the length of a geodesic representative. The following was stated by Thurston and proved by Brock.
Theorem 2.6 (Brock [Bro00]).
The length function $(\rho,\lambda)\mapsto\ell_{\rho}(\lambda)$ is continuous on the open subset of ${\mathcal{AH}}(S)\times{\mathcal{ML}}(S)$ consisting of pairs $(\rho,\lambda)$ such that $\rho\in{\mathcal{R}}(|\lambda|)$.
2.7. End structures
Let $M$ be an hyperbolic three-manifold with incompressible boundary, and let $C(M)$ be its convex core. See [Mar16, §3.12,5.5] for a definition of the relative ends of $M$, which describe an ideal boundary of $M$. A relative end $E$ corresponds to a decreasing nested sequence $V_{1}\supset V_{2}\supset\dots$ of connected open subsets of $M$ “going to infinity”, each homeomorphic to $S_{E}\times(0,1)$, where $S_{E}$ is a finite type oriented surface (possibly with cusps) that only depends on $E$ (and actually realized as a pleated surface inside $M$).
A relative end is either contained in a the complement of the convex core $C(M)$, in which case it is geometrically finite and it corresponds to a component of $\partial_{\infty}M$, or it is contained in $C(M)$, it is then geometrically infinite. Geometrically finite ends do not contain any closed geodesic, since all closed geodesics are contained in the convex core $C(M)$. On the other hand, a geometrically infinite end $E$, with associated surface $S_{E}$, is characterized by the existence of a sequence of simple closed geodesics $(\gamma^{*}_{n})_{n}$ in $S_{E}$ whose geodesic representatives $(\gamma_{n})_{n}$ in $M$ have bounded length and exit the end $E$. The limit of $(\gamma_{n}^{*})_{n}$ is a well-defined measurable geodesic lamination $\Lambda(E)\subset S_{E}$, called the ending lamination of $E$ [Thu97, Bon86]. It is an arational measurable geodesic lamination on $S_{E}$.
The end invariant of a Kleinian surface group $\rho\in{\mathcal{AH}}_{o}(S)$ can be encoded into a quadruplet $\operatorname{E}(\rho)=(P,\Sigma,m,L)$ where the parabolic locus $P$ is a simple essential multicurve on $S^{-}\sqcup S^{+}$, the moderate part $\Sigma$ is a union of connected components of $(S^{-}\sqcup S^{+})-P$, $m\in{\mathcal{T}}(\Sigma)$ and $L$ is a union of arational measured laminations in each component of $S(L)\coloneqq(S^{-}\sqcup S^{+})-(P\cup\Sigma)$. For a connected component $L^{\prime}$ of $L$, let $S(L^{\prime})$ denote the connected component of $S(L)$ containing $L^{\prime}$. In particular the end invariant $\operatorname{E}(\rho)$ decomposes into a topological end data $e=(P,L)$ and a conformal end data $c=(\Sigma,m)$. We also identify $e$ with the measurable geodesic lamination $P\cup L$ and denote by $S(e)$ the open subsurface of $S$ obtained as the union of $S(L)$ and colar neighborhoods of components of $P$.
The end structure $\operatorname{E}(\rho)$ is always doubly incompressible in the sense of [AL13, §2.9], i.e. $P\cup L$ has no parallel components. Any such end structure is realizable as the end invariant of a Kleinian surface group [Ohs11, NS12].
For a one-sided degenerated structure $\rho\in{\mathcal{AH}}_{o}^{+}(S)$, its parabolic locus $P$ is entirely contained in $S^{-}$ and its moderate part has the form $\Sigma^{-}\sqcup S^{+}\subset S^{-}\sqcup S^{+}$.
Baba and Ohshika described the combination of ending laminations, parabolic locus and bending laminations that occur for Kleinian surface groups [BO23]. It follows from their result that, for $\rho\in{\mathcal{AH}}_{o}^{+}(S;e^{-})$ where $e^{-}=P\cup L$, the bending lamination $\operatorname{b}^{+}(\rho)$ of the top boundary of its convex core fills $S(e^{-})$ with $e^{-}$, i.e. any essential closed curve on $S(e^{-})$ intersects $P\cup L$ or $\operatorname{b}^{+}(\rho)$ (in other words, any essential closed curve on $S$ intersects $P\cup L\cup\Sigma^{-}\cup\operatorname{b}^{+}(\rho)$). Hence it lies in the subspace <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ consisting of those measured laminations without closed leaves of weight $\geq\pi$, and which fill $S(e^{-})$ with $e^{-}$.
2.8. Metrics converging to an end structure
[…]
  1. (1)
    The metrics $m_{n}|_{\Sigma}$ converge to $m$ in the length spectrum, i.e. $\ell_{m_{n}}(\partial\overline{\Sigma})\to 0$ and $\ell_{m_{n}}(c)\to\ell_{m}(c)$ for all non-peripheral closed curve $c\subset\Sigma$.
  2. (2)
    For each component $L^{\prime}$ of $L$, the restrictions $m_{n}|_{S(L^{\prime})}$ converge to $L^{\prime}$, i.e. there is a non-peripheral simple closed curve $c\subset S(L^{\prime})$ such that $\ell_{m_{n}}(\partial\overline{S(L^{\prime})})/\ell_{m_{n}}(c)\to 0$ and any subsequence of $m_{n}|_{S(L^{\prime})}$ contains a subsequence converging to a projective measured lamination supported by $L^{\prime}$.
[…]
The following is a special case of Theorem D in [AL13]. The injectivity part is a consequence of the Brock-Canary-Minsky ending lamination theorem [BCM12], while the continuity and properness parts are due to Anderson and Lecuire [AL13]. The only part that we need is the second sentence.
Proposition 2.9.
If ${\mathcal{SH}}_{o}^{+}(S)$ denotes ${\mathcal{AH}}_{o}^{+}(S)$ with the strong topology instead of the algebraic topology, then the map $\operatorname{E}\colon{\mathcal{SH}}_{o}^{+}(S)\to{\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})$ which associates to $\rho$ its end structure is a homeomorphism, with inverse denoted $\overline{U}$.
In particular, if a sequence $(\rho_{n})_{n\in{\mathbb{N}}}\subset{\mathcal{QF}}_{o}(S)$ has conformal boundary $(m^{-}_{n},m^{+}_{n})_{n\in{\mathbb{N}}}\in{\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})$ converging to the end structure $\operatorname{E}(\rho_{\infty})=({\mathcal{E}}^{-},m^{+})\in{\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})$ of $\rho_{\infty}\in{\mathcal{AH}}_{o}^{+}(S)$, then $\rho_{n}\to\rho_{\infty}$ algebraically.
2.10. Quasi-conformal deformations of Kleinian groups
Let $\rho_{0}\in{\mathcal{AH}}(S)$ be a Kleinian surface group. A quasi-conformal deformation of $\rho_{0}$ is a point $\rho\in{\mathcal{AH}}(S)$ for which there exists a quasi-conformal homeomorphism $w\colon S^{2}\to S^{2}$ such that $w\circ\rho_{0}\circ w^{-1}=\rho$. Let ${\mathcal{QH}}(\rho_{0})\subset{\mathcal{AH}}(S)$ denote the subspace of quasi-conformal deformations of $\rho_{0}$. As defined this way, ${\mathcal{QH}}(\rho_{0})$ is still endowed with the algebraic topology. However, there is another natural topology on ${\mathcal{QH}}(\rho_{0})$, where two quasi-conformal deformations are close to each other if they are conjugate by a quasi-conformal homeomorphism with quasi-conformal constant close to $1$. Those two topologies actually coincide : for quasi-fuchsian groups it follows from Marden quasi-conformal stability theorem [Mar16, §5.1.2], in this case ${\mathcal{QH}}(\rho_{0})={\mathcal{QF}}_{o}(S)$ is the interior of ${\mathcal{AH}}_{o}(S)$ (if $\rho_{0}\in{\mathcal{AH}}_{o}(S)$), and in general from Matsuzaki conditional stability theorem [MT98, §7.4.2].
Combined with the simultaneous uniformization theorem, we have the following, see Theorems 7.53 and 7.55 in [MT98, §7.4.2].
Theorem 2.11.
Let $\rho_{0}\in{\mathcal{AH}}_{o}(S)$ with nontempy conformal boundary. Then simultaneous uniformization provides a biholomorphic homeomorphism
$U\colon{\mathcal{T}}(\partial_{\infty}\rho_{0})\overset{\cong}{\longrightarrow}{\mathcal{QH}}(\rho_{0}).$
Moreover, ${\mathcal{QH}}(\rho_{0})$ is a complex (regular) submanifold of an open subset of the representation variety of $\pi_{1}(S)$ into $PSL(2,{\mathbb{C}})$.
Let $e_{0}=(P_{0},L_{0})$ be the topological end invariants of $\rho_{0}\in{\mathcal{AH}}_{o}(S)$, i.e. $P_{0}$ is its parabolic locus and $L_{0}$ its ending lamination. Denote by ${\mathcal{AH}}_{o}(S;e_{0})$ the subspace of ${\mathcal{AH}}_{o}(S)$ consisting of those representations with the same topological end invariants. Then the ending lamination theorem [BCM12] gives the following.
Theorem 2.12 (Ending lamination theorem, Brock-Minsky-Canary [BCM12]).
Let $\rho_{0}\in{\mathcal{AH}}_{o}(S)$ with topological end invariant $e_{0}$. Then ${\mathcal{AH}}_{o}(S;e_{0})={\mathcal{QH}}(\rho_{0})$.
3. Continuity
3.1. Holonomy of projective structures
Let ${\mathcal{X}}(S)$ denote the character variety of $S$, i.e. the space of $\rm{PSL}(2,{\mathbb{C}})$-conjugacy classes of homomorphisms $\pi_{1}(S)\to\rm{PSL}(2,{\mathbb{C}})$. Let also ${\mathcal{P}}(S)$ be the space of complex projective structures on $S$, i.e. $(\rm{PSL}(2,{\mathbb{C}}),{\mathbb{CP}}^{1})$-structures on $S$.
To a projective structure $Z$ on $S$ one can associate its developing map $\operatorname{dev}_{Z}\colon\tilde{S}\to{\mathbb{CP}}^{1}$, which is an orientation-preserving local homeomorphism, and its holonomy representation $\operatorname{hol}_{Z}\colon\pi_{1}(S)\to\rm{PSL}(2,{\mathbb{C}})$. This defines a map
(1) $\operatorname{hol}\colon{\mathcal{P}}(S)\to{\mathcal{X}}(S)$
which is a local biholomorphic homeomorphism [Dum09, Theorem 5.1].
Denote by ${\mathcal{P}}_{0}(S)$ the subspace of projective structures with injective developing map (it is closed in ${\mathcal{P}}(S)$ [KT92, Prop. 3.1.2]). The holonomy representation $\operatorname{hol}_{Z}$ of such a projective structure $Z\in{\mathcal{P}}_{0}(S)$ is actually discrete and faitfhul, and its domain of discontinuity $\Omega(\operatorname{hol}_{Z})\subset S^{2}$ has an invariant simply-connected component : the image of $\tilde{S}$ under $\operatorname{dev}_{Z}$. Denote by ${\mathcal{B}}(S)$ the subspace of ${\mathcal{X}}(S)$ consisting of representations with this property (i.e. discrete, faithful, with a simply-connected invariant component of the domain of discontinuity). The map $\operatorname{hol}$ restricts to a biholomorphic homeomorphism [KT92, Prop. 3.2.3]
(2) $\operatorname{hol}|_{{\mathcal{P}}_{0}(S)}\colon{\mathcal{P}}_{0}(S)\to{\mathcal{B}}(S).$
3.2. Thurston’s grafting parameterization of projective structures
We want to prove that the map $\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S)\to{\mathcal{ML}}(S^{+})$ sending a one-sided degenerated structure to the bending lamination of the top boundary of its convex core is continuous. This directly follows from Thurston’s grafting parameterization of complex projective structures on $S$, as is detailed in this section and the next one.
The space ${\mathcal{P}}(S)$ of projective structures on $S$ admits a geometric parameterization in terms of the grafting construction. Given a hyperbolic metric $m\in{\mathcal{T}}(S)$ and a measured lamination $\lambda\in{\mathcal{ML}}(S)$, one can construct a complex projective structure $\operatorname{Gr}_{\lambda}(m)$ called the grafting of $m$ along $\lambda$. When $\lambda$ is a weighted multicurve, the grafting is obtained by cutting $(S,m)$ along the curves of $\lambda$ and gluing in complex cylinders of lengths given by the weights of $\lambda$. This construction extends by continuity to all measured laminations. This describes Thurston’s grafting parameterization
(3) $\operatorname{Gr}\colon{\mathcal{T}}(S)\times{\mathcal{ML}}(S)\to{\mathcal{P}}(S)$
which is a homeomorphism [KT92].
The inverse map $\operatorname{Gr}^{-1}$, when restricted to the space ${\mathcal{P}}_{0}(S)$ of projective structures with injective developing map, can be described geometrically through convex pleated surfaces. Given a projective structure $Z\in{\mathcal{P}}_{0}(S)$, its holonomy $\operatorname{hol}_{Z}$ stabilizes the component $\Omega\coloneqq\operatorname{dev}_{Z}(\tilde{S})\subset{\mathbb{CP}}^{1}$ of its domain of discontinuity. Seeing ${\mathbb{CP}}^{1}$ as the conformal boundary at infinity of ${\mathbb{H}}^{3}$, denote by $\operatorname{Pl}(Z)$ the boundary of the convex hull of ${\mathbb{CP}}^{1}-\Omega$ inside ${\mathbb{H}}^{3}$. It is a convex pleated plane, and it is $\pi_{1}(S)$-invariant by construction. With the induced path metric, $\operatorname{Pl}(Z)$ is isometric to ${\mathbb{H}}^{2}$, hence its quotient by the $\pi_{1}(S)$-action defines a point in ${\mathcal{T}}(S)$. Similarly, the bending measured lamination of $\operatorname{Pl}(Z)$ defines a point in ${\mathcal{ML}}(S)$. This pair of geometric data coincides with $\operatorname{Gr}^{-1}(Z)$.
3.3. One-sided degenerated structures
The space of one-sided degenerated structures ${\mathcal{AH}}_{o}^{+}(S)$ is a subspace of ${\mathcal{AH}}(S)\subset{\mathcal{X}}(S)$. Actually, it is a subspace of ${\mathcal{B}}(S)$. Indeed, consider a one-sided degenerated structure $\rho\in{\mathcal{AH}}_{o}^{+}(S)$. By definition, the top end of $M_{\rho}$ is non-degenerated, hence the corresponding component of the domain of dicontinuity $\Omega^{+}(\rho)\subseteq\Omega(\rho)$ is simply-connected and invariant under $\rho$. Therefore, $\rho$ lies in ${\mathcal{B}}(S)$ and under (2) corresponds to a projective structure $Z_{\rho}\coloneqq\operatorname{hol}|_{{\mathcal{P}}_{0}(S)}^{-1}(\rho)\in{\mathcal{P}}_{0}(S)$ with injective developing map $\operatorname{dev}_{Z_{\rho}}$. Moreover, the image of $\operatorname{dev}_{Z_{\rho}}$ coincides with $\Omega^{+}(X)$. In particular, the convex pleated surface $\operatorname{Pl}(Z_{X})$ coincides with the top side of the convex core of $X$, hence their bending laminations coincide.
All in all, this identifies the bending map $\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S)\to{\mathcal{ML}}(S^{+})$ with the continuous map
$\operatorname{proj}_{{\mathcal{ML}}(S)}\circ\operatorname{Gr}^{-1}\circ\operatorname{hol}|_{{\mathcal{P}}_{0}(S)}^{-1}.$
Proposition 3.4.
The bending map $\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S)\to{\mathcal{ML}}(S^{+})$ is continuous on the space of one-sided degenerated structures on $S\times{\mathbb{R}}$.
Corollary 3.5.
The map $\operatorname{B}^{+}\colon[0,\infty]\times{\mathcal{T}}(S^{+})\to{\mathcal{ML}}(S^{+})$ defined in §1.1 is continuous.
Proof.
By the above Proposition, $\operatorname{b}^{+}$ is continuous. The map $\overline{U}$ is continuous by Proposition 2.9, while $\overline{r}$ is continuous by construction. Thus $\operatorname{B}^{+}$ is continuous. ∎
4. Properness
The goal of this section is to prove that the map $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;e^{-})\to{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper (see the end of Section 2.7 for a description of the bending laminations that occur for strucutes in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$). This will imply that $\operatorname{b}^{+}_{\infty}\colon{\mathcal{T}}(S^{+})\to{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper too. The proof is similar to the proof of properness of similar maps by Baba and Ohshika in [BO23] and Anderson and Lecuire in [AL13].
Proposition 4.1.
Let $e^{-}=(P,L)$ be a topological end structure on $S^{-}$. Let $\Sigma^{-}\coloneqq S^{-}-(P\cup S(L))$ be the moderate subsurface of the bottom end, and let $S^{+}=S$ be the one of the top end.
Let <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ be the subspace of ${\mathcal{ML}}(S^{+})$ consisting of the measured laminations $\mu$ satisfying the following conditions :
  1. (1)
    closed leaves of $\mu$ have weight strictly smaller than $\pi$ ;
  2. (2)
    $\mu$ and $e^{-}$ together fill up the immoderate surface $S(e^{-})$ (identifying $S(e^{-})$ with a subsurface of $S^{+}$), i.e. for any essential closed curve $c$ on $S(e^{-})$ one has $i(c,\mu)+i(c,e^{-})>0$ (where in the last inequality one needs to pick a measured lamination representing $e^{-}$).
Consider the map
$\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;e^{-})\longrightarrow{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$
that associates to a structure $\rho$ the pair $(\operatorname{m}^{-}(\rho),\operatorname{b}^{+}(\rho))$ of the conformal structure at infinity of the moderate surface of the bottom end and the bending lamination of the top side of the convex core.
Then $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is proper.
Proof.
Consider a sequence of representations $(\rho_{n})_{n}$ in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ with convergent (b)ending invariants $(\operatorname{m}^{-}(\rho_{n}),\operatorname{b}^{+}(\rho_{n}))_{n}\subset{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$. By Proposition 4.3, it has an algebraically convergent subsequence. By Corollary 4.8, this algebraic limit belongs to ${\mathcal{AH}}_{o}(S;e^{-})$. This concludes the proof. ∎
4.2. Algebraic convergence
The first step of the proof of properness is the following.
Proposition 4.3.
Let $(\rho_{n})_{n}$ be a sequence of representations in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ such that $(\operatorname{m}^{-}\times\operatorname{b}^{+}(\rho_{n}))_{n}$ converges in ${\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$. Then $(\rho_{n})_{n}$ has an algebraically convergent subsequence.
Proof.
Write $(\operatorname{m}^{-}\times\operatorname{b}^{+})(\rho_{n})=(m_{n}^{-},\lambda^{+}_{n})$ for each $n\in{\mathbb{N}}$, and $(m^{-},\lambda^{+})$ their limit.
Let us assume that $(\rho_{n})_{n}$ has no algebraically convergent subsequence. By Morgan-Shalen theorem 2.3 [MS84] combined with Skora’s description of geometric actions on trees [Sko96] (see 2.2.12 and 2.3.5 in [Ota96]), up to extracting a subsequence, $(\rho_{n})_{n}$ converges to the action of $\Gamma=\pi_{1}(S)$ on the real tree ${\mathcal{T}}_{\mu}$ dual to a measured lamination $\mu\in{\mathcal{ML}}(S)$, i.e. there is a sequence of positive real numbers $\epsilon_{n}\searrow 0$ such that $\epsilon_{n}\ell_{\rho_{n}}(\gamma)\to i(\mu,\gamma)$ for all $\gamma\in\Gamma$ (where $\ell_{\rho_{n}}(\gamma)$ denotes the length of a geodesic representative of $\gamma$ in $M_{\rho_{n}}\coloneqq{\mathbb{H}}^{3}/\rho_{n}(\Gamma)$, and is zero if $\rho_{n}(\gamma)$ is parabolic).
We will show that $\mu$ cannot intersect $\Sigma^{-}$ nor $e^{-}=P\cup L$ nor $\lambda^{+}$, which then contradicts the condition that $e^{-}\cup\lambda^{+}$ must fill $S(e^{-})$. Note that this argument also applies in the case where $\lambda^{+}=0$, since in that case $\Sigma^{-}=S^{+}$ and $e^{-}=\emptyset$.
Since $\epsilon_{n}\ell_{\rho_{n}}(c)\to i(\mu,c)$ for any essential closed curve $c$ on $S$ one has the implication
(4) $i(\mu,c)>0\quad\Longrightarrow\quad\ell_{\rho_{n}}(c)\to\infty.$
First, since $P$ consists of parabolics in $M_{\rho_{n}}$, we have $\ell_{\rho_{n}}(P)=0$ for each $n$, hence $i(\mu,P)=0$ by (4). Thus, each component of $\mu$ is either contained in $\Sigma^{-}$ or in $S(L)$. Suppose that $\mu$ has a component contained in $\Sigma^{-}$, then there is an essential closed curve $c$ in $\Sigma^{-}$ with $i(\mu,c)>0$, hence $(\ell_{\rho_{n}}(c))_{n}$ is unbounded by (4). But $(\ell_{m^{-}_{n}}(c))_{n}$ is bounded since $(m^{-}_{n})_{n}$ is bounded in ${\mathcal{T}}(\Sigma^{-})$. By Sullivan’s convex hull theorem [EM87], $(\Sigma^{-},m^{-}_{n})$ is $K$-quasi-isometric to the boundary of the convex core of $M_{\rho_{n}}$ facing it, with $K$ a universal constant. In particular, this implies that $(\ell_{\rho_{n}}(c))_{n}$ is bounded, a contradiction. Therefore, $\mu$ has support contained in $S(L)=S^{-}-(P\cup\Sigma^{-})$.
In [AL13], Anderson and Lecuire show that no component of $L$ can be realized in the dual tree ${\mathcal{T}}_{\mu}$ (see Section 2.2 above, or Claim 3.5 and Lemma 3.6 in [AL13]). In particular, by Proposition 3.1.1 in [Ota96], we obtain that no component of $L$ intersects $\mu$ (in the sense of §2.2, see [Ota96, §3]). But $\mu\subset S(L)$, and each component of $L$ is arational (i.e. minimal, maximal and irrational) in its respective component of $S(L)$, hence the support of $\mu$ must be contained in $L$ [Ota96, A.3.10].
Finally, Bridgeman proved that there is a constant bounding the length of bending laminations of convex cores [Bri98, Prop. 2] (see §2.4 for the definition of the length of a measured lamination in a Kleinian surface manifold). But since $e^{-}$ and $\lambda^{+}$ fill $S(e^{-})$ and $\mu$ has support contained in $L\subset e^{-}$, we obtain that $i(\mu,\lambda^{+})>0$. We claim that this forces $(\ell_{\rho_{n}}(\lambda^{+}))_{n}$ to be unbounded, a contradiction with Bridgeman’s result.
Since $i(\mu,\lambda^{+})>0$, there is a minimal component $\lambda^{+}_{0}\subset\lambda^{+}$ with $i(\mu,\lambda^{+}_{0})>0$. If $\lambda^{+}_{0}$ is a closed curve, then (4) implies that $(\ell_{\rho_{n}}(\lambda^{+}_{0}))_{n}$ is unbounded, so is $(\ell_{\rho_{n}}(\lambda^{+}))_{n}$. Let us now assume that $\lambda^{+}_{0}$ is irrational. We cannot apply (4) to prove the claim since that implication is a priori only valid for closed curves (and, by linearity, for weighted multicurves). Better control on (4) is given by Otal’s continuity theorem [Ota96, Theorem 4.0.1], of which we need the more general form stated by Lecuire in [Lec20, Theorem 3.7].
Theorem 4.4 (Otal continuity theorem).
Let $(\rho_{n})_{n\in{\mathbb{N}}}\subset{\mathcal{AH}}(S)$ be a sequence converging to a small minimal action of $\pi_{1}(S)$ on a real tree ${\mathcal{T}}_{\mu}$, where $\mu\in{\mathcal{ML}}(S)$. Let $\epsilon_{n}\searrow 0$ be such that $\epsilon\ell_{\rho_{n}}\to i(\mu,-)$. Let $s_{0}\in{\mathcal{T}}(S)$ be any hyperbolic metric on $S$. Let $\alpha\subset S$ be a geodesic lamination realized in ${\mathcal{T}}_{\mu}$, i.e. each leaf of $\alpha$ intersects $\mu$ transversely. Then there exists a neighborhood ${\mathcal{V}}(\alpha)\subset{\mathcal{GL}}(S)$ of $\alpha$ and constants $K$, $n_{0}$ such that, for any simple closed curve $c\subset{\mathcal{V}}(\mu)$ and for any $n\geq n_{0}$,
(5) $\epsilon_{n}\ell_{\rho_{n}}(c)\geq K\ell_{s_{0}}(c).$
Pick a sequence of weighted simple closed curves $(\alpha_{j})_{j\in{\mathbb{N}}}$ converging to $\lambda^{+}_{0}$ in ${\mathcal{ML}}(S)$. By Otal’s continuity theorem there exist constants $K$, $N$ and $J$ such that, for all $n\geq N$ and $j\geq J$,
$\ell_{\rho_{n}}(\alpha_{j})\geq\frac{K}{\epsilon_{n}}\ell_{s_{0}}(\alpha_{j}),$
where $s_{0}$ is any fixed reference metric, independent of $n$ and $j$. By continuity of length functions (Theorem 2.6 by Brock [Bro00]), the limit $j\to\infty$ gives, for each $n\geq N$,
$\ell_{\rho_{n}}(\lambda^{+}_{0})\geq\frac{K}{\epsilon_{n}}\ell_{s_{0}}(\lambda^{+}_{0}).$
Since $\epsilon_{n}\searrow 0$, while $K\ell_{s_{0}}(\lambda^{+}_{0})$ is a positive constant, we obtain that $\ell_{\rho_{n}}(\lambda^{+}_{0})\to\infty$, concluding the proof. ∎
4.5. Geometric convergence
Let $(\rho_{n})_{n}$ be a sequence of representations in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ as in Proposition 4.3 with $\operatorname{m}^{-}\times\operatorname{b}^{+}(\rho_{n})=(m^{-}_{n},\lambda^{+}_{n})$ converging to $(m^{-},\lambda^{+})$ in ${\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$. By Proposition 4.3, up to replacing $(\rho_{n})_{n}$ by a subsequence, we assume that it converges algebraically to a representation $\rho_{\infty}\in{\mathcal{AH}}_{o}(S)$. The goal of this section is to show that $\rho_{\infty}$ belongs to ${\mathcal{AH}}_{o}^{+}(S;e^{-})$, i.e. that ${\mathbb{H}}^{3}/\rho_{\infty}(\Gamma)$ has parabolic locus $P$, no new geometrically infinite end, and that it still has $L$ as its ending lamination.
First, the bottom end is controlled.
Lemma 4.6 (Anderson-Lecuire).
The bottom end of the algebraic limit ${\mathbb{H}}^{3}/\rho_{\infty}(\Gamma)$ is geometrically finite on the moderate subsurface $\Sigma^{-}$, without new rank $1$ cusps on $\Sigma^{-}$, and has ending lamination $L$ in the immoderate subsurface $S^{-}-(P\cup\Sigma^{-})$.
Proof.
The moderate part of the statement is [AL13, Lemma 4.1] and the immoderate part is [AL13, Lemma 4.2]. ∎
The top end can be controlled thanks to the following result of Bonahon-Otal [BO04, §5,6] and Lecuire [Lec04, §4.1], [Lec06, §4.2]. We state a slightly weaker version, although their result also works in presence of parabolics in the top end.
Proposition 4.7 (Bonahon-Otal, Lecuire).
Let $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ be a sequence of convex pleated surfaces such that $(\rho_{n})_{n}$ converges algebraically to some $\rho_{\infty}$. Let $(\lambda_{n})_{n}$ be the sequence of bending laminations of those pleated surfaces. Assume that $\lambda_{n}\to\lambda$ in ${\mathcal{ML}}(S)$ and that the closed leaves of $\lambda_{n}$ and $\lambda$ have weight $<\pi$. Then $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ converges geometrically to a convex pleated surface $(f_{\infty},\pi_{1}(S),\rho_{\infty})$ inside ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$. In particular, the top boundary of the convex core of ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$ is homeomorphic to $S^{+}$.
Thus, we obtain the following, which concludes the proof that $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is proper.
Corollary 4.8.
The algebraic limit ${\mathbb{H}}^{3}/\rho_{\infty}(\Gamma)$ belongs to ${\mathcal{AH}}_{o}^{+}(S;e^{-})$.
Actually, since ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$ has no new parabolics, this also implies that the sequence $(\rho_{n})$ converges strongly to $\rho_{\infty}$ (by Anderson-Canary, see [Mar16, Theorem 4.6.2]), although we do not need this.
5. Real-analyticity
Lemma 5.1.
Let $e^{-}=P\cup L$ be a topological end structure on $S^{-}$ and $(m^{-},\lambda^{+})\in{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$. Then $(\operatorname{m}^{-}\times\operatorname{b}^{+})^{-1}(m^{-},\lambda^{+})$ is a real-analytic subset of ${\mathcal{AH}}_{o}^{+}(S;e^{-})$.
Proof.
The fibre $F\coloneqq(\operatorname{m}^{-}\times\operatorname{b}^{+})^{-1}(m^{-},\lambda^{+})$ is the intersection of $(\operatorname{m}^{-})^{-1}(m^{-})$ and $(\operatorname{b}^{+})^{-1}(\lambda^{+})$, hence it suffices to show that each of these two subsets of ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ is real-analytic.
First, Theorems 2.11 and 2.12 say that ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ is a complex regular submanifold of an open subset of ${\mathcal{X}}(S)$, biholomorphically homeomorphic to ${\mathcal{T}}(\Sigma^{-})\times{\mathcal{T}}(S^{+})$. In particular, $(\operatorname{m}^{-})^{-1}(m^{-})$ is a complex regular submanifold of an open subset of ${\mathcal{X}}(S)$, biholomorphically homeomorphic to ${\mathcal{T}}(S^{+})$.
Pick a maximal geodesic lamination $L$ containing the support of $\lambda^{+}$. By Bonahon’s shear-bend parameterization Theorem 2.5, the open subset ${\mathcal{R}}(L)\subset{\mathcal{X}}(S)$ consisting of those representations realizing a pleated surface with pleating locus contained in $\lambda^{+}$ is biholomorphically homeomorphic to an open subset of ${\mathcal{H}}(L;{\mathbb{C}}/2\pi i{\mathbb{Z}})$ via the shear-bend coordinates $\sigma+i\beta$. For a generic $\rho\in{\mathcal{R}}(L)$, the associated pleated surface $P_{\rho}$ lies inside the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ and is neither convex nor concave, in particular $\beta(\rho)$ takes positive and negative values in $[-\pi,\pi]$, i.e. it does not correspond to a transverse measure. However, if $\rho\in{\mathcal{R}}(L)$ does realize $P_{\rho}$ as the upper boundary of the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$, then $\beta(\rho)$ corresponds to a transverse measure with support contained in $L$. In particular, for $\rho_{0}\in{\mathcal{AH}}_{o}^{+}(S;e^{-})$ with $\operatorname{b}^{+}(\rho_{0})=\lambda^{+}$, the real-analytic subset $\beta^{-1}(\beta(\rho_{0}))\subset{\mathcal{R}}(L)$ consists of those representations realizing a pleated surface with the same bending cocycle as $\rho_{0}$, that is, in this case, a convex pleated surface with bending lamination $\lambda^{+}$. In other words, $(\operatorname{b}^{+})^{-1}(\lambda^{+})\cap{\mathcal{R}}(L)=\beta^{-1}(\beta(\rho_{0}))\cap{\mathcal{R}}(L)$. Since ${\mathcal{R}}(L)$ is open in ${\mathcal{X}}(S)$, this shows that $(\operatorname{b}^{+})^{-1}(\lambda^{+})$ is locally real-analytic around $\rho_{0}$. Thus it is real-analytic. ∎
In fact, ${\mathcal{AH}}_{o}^{+}(S;e^{-})\subset{\mathcal{R}}(L)$ as follows from the proof of the density theorem by Namazi-Souto [NS12] and Ohshika [Ohs11] : only the ending laminations are unrealizable.
For our purpose, the fact that the fibres of $\operatorname{m}^{-}\times\operatorname{b}^{+}$ are real-analytic is important for two reasons, recollected in the following two results.
Lemma 5.2.
Compact real-analytic spaces are absolute neighborhood retracts (see §6.6).
Proof.
Combine [Sul71] with [Dav86, Corollary 8A] or see [DS24, Lemma 4.8]. ∎
Theorem 5.3 (Borel-Haefliger [BH61], Sullivan [Sul71]).
A compact real-analytic space $X$ has a fundamental class modulo two, i.e. a nonzero homology class in $H_{\dim(X)}(X;{\mathbb{Z}}/2)$.
6. Injectivity
As mentioned in the introduction, a key ingredient in showing the main result is to first show the following partial result, where the parabolic locus and ending lamination are empty.
Theorem 6.1.
Consider the map $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{ML}}_{<\pi}(S^{+})$ sending a quasifuchsian representation $\rho$ to the pair $(\operatorname{m}^{-}(\rho),\operatorname{b}^{+}(\rho))$ of the conformal structure at infinity of the bottom end and the pleating measured lamination of the top boundary of the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$. It is a homeomorphism.
The proof is very similar to the proof of the injectivity of $\operatorname{b}^{-}\times\operatorname{b}^{+}$ [DS24]. After proving it, we conclude the proof of the main result.
6.2. Prescribed third fundamental form
Given $\rho\in{\mathcal{QF}}_{o}(S)$, a theorem of Labourie [Lab92] states that each end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ is uniquely foliated by smooth closed subsurfaces of constant curvature equal to $K$, with $K$ contained in the interval $(-1,0)$. The third fundamental form of such a subsurface has constant curvature $K^{*}\coloneqq K/(K+1)$. Thus it gives a point in the rescaled Teichmüller space ${\mathcal{T}}^{K^{*}}(S)$, which is the space of isotopy classes of metrics of constant curvature $K^{*}$ on $S$. Focusing on the top end, it follows that, for each $K\in(-1,0)$, one can define the map
(6) $III^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}^{K^{*}}(S)$
sending a quasifuchsian representation $\rho$ to the third fundamental form of the constant curvature $K$ surface in the top end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$. The third fundamental form $III^{+}_{K}$ is a “smooth measure of bending” and Chen-Schlenker proved that the analogous “smooth version” of Proposition 6.1 holds (see Corollary 1.4 and the proof of Lemma 4.1 in [CS25]). We need the following special case of their result, which is more general.
Theorem 6.3 (Chen-Schlenker [CS25]).
The map $\operatorname{m}^{-}\times III^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{T}}^{K^{*}}(S^{+})$ is a homeomorphism.
This is a mixed version of the earlier result of Schlenker that $III^{-}_{K}\times III^{+}_{K}$ is a homeomorphism [Sch06], which was an essential ingredient in [DS24].
These smooth analogues of $\operatorname{m}^{-}\times\operatorname{b}^{+}$ actually approximate $\operatorname{m}^{-}\times\operatorname{b}^{+}$, as follows from the following result of Bonsante-Mondello-Schlenker [BMS13] and Belraouti [Bel17]. See the version stated in [DS24].
Proposition 6.4.
When $K\to-1$, the maps $III^{+}_{K}$ converge to $\operatorname{b}^{+}$ in the length spectrum, uniformly on compacts. More precisely, $\ell_{III^{+}_{K}(\rho)}(c)\to i(\lambda^{+}(\rho),c)$ when $K\to-1$, for each closed curve $c$ in $S$ and each $\rho\in{\mathcal{QF}}_{o}(S)$, and the convergence is uniform for $\rho$ in a compact subset of ${\mathcal{QF}}_{o}(S)$.
In order to apply the variant of Finney’s theorem as in the next section, we identify the targets of the maps $III^{+}_{K}$, for $K\in(-1,0)$, and $\operatorname{b}^{+}$, via the earthquake map $E\colon{\mathcal{ML}}(S)\times{\mathcal{T}}(S)\to{\mathcal{T}}(S)\colon(\lambda,m)\mapsto E(\lambda,m)$ as in [DS24, §4.1]. More precisely, we fix a point $m_{0}\in{\mathcal{T}}(S)$ and consider, for each $K\in(-1,0)$, the map
(7) $V_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{+})\colon\rho\mapsto\frac{1}{\sqrt{\left|K^{*}\right|}}E(-,m_{0})^{-1}(\left|K^{*}\right|III^{+}_{K}(\rho)),$
which is such that $\operatorname{m}^{-}\times V_{K}$ is a homeomorphism since the earthquake map is a homeomorphism by Thurston’s earthquake theorem [Thu86]. Using the estimate [DS24, Lemma 4.4], we obtain the following.
Corollary 6.5 ([DS24, Lemma 4.2]).
The maps $V_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{+})$ converge to $\operatorname{b}^{+}$ uniformly on compacts, when $K\to-1$ (with respect to the induced metric when embedding ${\mathcal{ML}}(S^{+})$ inside ${\mathbb{R}}^{\{c_{1},\dots,c_{N}\}}$ for a large enough collection of closed curves).
6.6. Fibres of a limit of homeomorphism
Let $\operatorname{\mathbb{S}}^{k}\subset{\mathbb{D}}^{k+1}$ denote the inclusion of the $k$-sphere as the boundary of the $(k+1)$-ball. Recall that a space $K$ is weakly contractible if, for any $k\geq 0$, any map $\operatorname{\mathbb{S}}^{k}\to K$ extends to a map ${\mathbb{D}}^{k+1}\to K$. If a space $X$ is weakly contractible, then it is acyclic, i.e. all its reduced homology groups are trivial.
A subset $K\subset N$ is a neighbourhood retract if it has an open neighbourhood $\Omega\subset N$ such that the inclusion $j\colon K\xhookrightarrow{}\Omega$ has a retraction $r\colon\Omega\to K$, i.e. $r\circ j=\operatorname{id}_{K}$. A metrisable space $X$ is an absolute neighbourhood retract (ANR) if, for any closed embedding $X\xhookrightarrow{}Y$ into a metrisable space $Y$, $X$ is a neighbourhood retract in $Y$.
A sequence of maps $(f_{n}\colon x\to Y)_{n\in{\mathbb{N}}}$ converges continuously to $f_{\infty}$ if $f_{n}(x_{n})\to f(x)$ whenever $x_{n}\to x$. When $Y$ is metrisable and $f_{\infty}$ is continuous, continuous convergence coincides with compact convergence, i.e. uniform convergence on compact subsets of $X$ [Rem91, p. 98].
The following result is a variant of the theorem of Finney [Fin67] used in [DS24, Theorem 4.6].
Theorem 6.7.
Let $N$ and $N^{\prime}$ be two manifolds of the same dimension. Let $(f_{n}\colon N\to N^{\prime})_{n\in{\mathbb{N}}}$ be a sequence of open embeddings converging continuously (equivalently, compactly) to the continuous map $f_{\infty}\colon N\to N^{\prime}$, i.e. $f_{n}(x_{n})\to f_{\infty}(x_{\infty})$ whenever $x_{n}\to x_{\infty}$. If a fibre of $f_{\infty}$ is compact and is an absolute neighbourhood retract, then it is weakly contractible.
In fact, a compact weakly contractible ANR is automatically contractible, although we only need the fact that it is acyclic.
Proof.
Let $K\coloneqq f_{\infty}^{-1}(\ast)$ be a compact fibre, which is a retract of an open neighbourhood $\Omega$, which we can assume to have compact closure. Assuming that $K$ is not weakly contractible, we will show that $f_{\infty}$ sends a point of $\partial\Omega$ to $\ast$.
By contradiction, assume that there exists $k\geq 0$ and a map $\sigma\colon S^{k}\to K$, which is not contractible, i.e. does not extend to a map ${\mathbb{D}}^{k+1}\to K$. In particular, there is also no map ${\mathbb{D}}^{k+1}\to\Omega$ extending $\sigma$, for one could compose it with a retraction $\Omega\to K$ to obtain a map which lands in $K$.
Let $U\subset N^{\prime}$ be a contractible neighbourhood of $\ast$. Since $K$ is compact, we have $f_{n}(K)\subset U$ for all $n$ large enough. Fix such a $n\in{\mathbb{N}}$. Thus, $f_{n}(K)\subset U$. Since $U$ is contractible, the map $f_{n}\circ\sigma\colon\operatorname{\mathbb{S}}^{k}\to U$ extends to a map $g\colon{\mathbb{D}}^{k+1}\to U$. But recall that $f_{n}$ is an open embedding, hence $f_{n}\circ\sigma$ is not contractible inside $f_{n}(\Omega)$. Thus $g({\mathbb{D}}^{k+1})$ cannot be entirely contained in $f_{n}(\Omega)$, so it must intersect $f_{n}(\partial\Omega)$, say at a point $f_{n}(x_{U})$, for otherwise $g({\mathbb{D}}^{k+1})$ would be disconnected. In particular, $x_{U}\in\partial\Omega$ and $f_{n}(x_{U})\in U$.
Considering a nested sequence of contractible neighbourhoods $U_{0}\supset U_{1}\supset\dots\ni\ast$ such that $\bigcap_{i=0}^{\infty}U_{i}=\ast$, the above paragraph gives a sequence $(n_{i})_{i\in{\mathbb{N}}}$ in ${\mathbb{N}}$, with $n_{i}\to\infty$, and a sequence of points $(x_{i})_{i\in{\mathbb{N}}}$ in $\partial\Omega$, such that $f_{n_{i}}(x_{i})\in U_{i}$ for all $i\in{\mathbb{N}}$.
By compactness of $\partial\Omega$, up to extracting a subsequence, $(x_{i})_{i\in{\mathbb{N}}}$ converges to some $x_{\infty}\in\partial\Omega$. By continuous convergence, we obtain
$f_{\infty}(x_{\infty})=\lim_{i\to\infty}f_{n_{i}}(x_{i})=\ast,$
which is impossible since $\partial\Omega$ does not meet $K=f_{\infty}^{-1}(\ast)$. This concludes the proof that $K$ is weakly contractible. ∎
We are finally able to prove Proposition 6.1, i.e. that $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{ML}}_{<\pi}(S^{+})$ is a homeomorphism, as in [DS24] for $\operatorname{b}^{-}\times\operatorname{b}^{+}$.
Proof of Theorem 6.1.
First, using Proposition 4.1 with empty parabolic locus and ending lamination, the map $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is proper. In particular, its fibres are compact. Let $F$ be a fibre of $\operatorname{m}^{-}\times\operatorname{b}^{+}$.
By Lemmas 5.1 and 5.2, $F$ is a real-analytic space and an ANR.
Since $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is the limit of the homeomorphisms $\operatorname{m}^{-}\times V_{K}$ and the convergence is uniform on compacts, $F$ is weakly contractible by Theorem 6.7, hence it is acyclic.
Since $F$ is a compact real-analytic space, it has a fundamental class modulo two (by Borel-Haefliger and Sullivan, see Theorem 5.3), i.e. a nonzero homology class $\alpha\in H_{\dim(F)}(F;{\mathbb{Z}}/2)$. But $F$ is acyclic (in fact, contractible), hence its homology is trivial in all degrees except 00. Since $\alpha\neq 0$, we must have $\dim(F)=0$. Therefore, $F$ is a contractible compact real-analytic space of dimension 00, that is, a point. This proves that $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is injective. By invariance of domains, it is an open embedding. This concludes, since a proper open embedding into a connected space is a homeomorphism. ∎
6.8. Conclusion
We now have all the required ingredients to conclude the proof of the main result.
Proof of the main theorem.
Since
$\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;e^{-})\to{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$
is proper (Proposition 4.1), continuous (on the first factor, by the simultaneous uniformization theorem, see Section 2.10 ; on the second factor, by Proposition 3.4), and its domain and target are manifolds of the same dimension (By [Ker92, Theorem 4.7], <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is a manifold of the same dimension as ${\mathcal{T}}(S^{+})$), it suffices to show that it is injective. With the notation of Section 1.1, it suffices to show that $\operatorname{b}^{+}_{\infty}$ is injective. But it is the limit of the open embeddings $\operatorname{b}^{+}_{t}$, when $t\to\infty$, and the convergence is continuous since $\operatorname{B}^{+}$ is continuous. By following the proof of Proposition 6.1 and using Theorem 6.7, we obtain that $\operatorname{b}^{+}_{\infty}$ is injective, hence $\operatorname{b}^{+}$ is injective too. ∎

### v2

@@
 structures $\rho\in{\mathcal{AH}}_{o}(S)$ whose top end is not degenerated.
-Refer to caption$\partial_{\infty}^{+}M$$\partial_{\infty}^{-}M$$C(M)$$\cong S$$C(M)$$\partial^{+}C(M)$$\partial^{-}C(M)$cuspdegenerate end
+[…]
 Figure 1.
-Schematic representation of hyperbolic manifolds homeomorphic to $S\times{\mathbb{R}}$, with $1$-dimensional “$S$-direction”, conformal boundary in red and convex core in grey.
+Schematic representation of hyperbolic manifolds homeomorphic to $S\times{\mathbb{R}}$, with 11-dimensional “$S$-direction”, conformal boundary in red and convex core in grey.
 The structures are, from left to right: Fuchsian, quasi-Fuchsian, Kleinian with a cusp and Kleinian with simply-degenerated bottom end.
@@
 Denote by ${\mathcal{AH}}_{o}^{+}(S;P\cup L)\subset{\mathcal{AH}}_{o}(S)$ the subspace of one-sided degenerated structures with bottom end cusped along $P$ and having ending laminations $L$.
-Let <${\mathcal{ML}}^{\perp P\cup L}_{<\pi}(S^{+})\subset{\mathcal{ML}}(S^{+})$ be the subspace of measured laminations $\mu$ on $S^{+}$ without closed leaves of weight $\geq\pi$ and such that $\mu$ and $P\cup L$ fill $S-\Sigma^{-}$, i.e.
+Let ${\mathcal{ML}}^{\perp P\cup L}_{<\pi}(S^{+})\subset{\mathcal{ML}}(S^{+})$ be the subspace of measured laminations $\mu$ on $S^{+}$ without closed leaves of weight $\geq\pi$ and such that $\mu$ and $P\cup L$ fill $S-\Sigma^{-}$, i.e.
 any essential closed curve on $S$ intersects either $\mu\cup P\cup L$ or $\Sigma^{-}$.
@@
 Then, in section 4 we prove properness.
-The argument follows the same outline as for similar maps, considered by Bonahon-Otal [BO04], Lecuire [Lec06], Anderson-Lecuire [AL13] and Baba-Oshika [BO23].
-We first show that any sequence of representations with convergent image under $\operatorname{m}^{-}\times\operatorname{b}^{+}$ must have an algebraically convergent subsequence, using the theory of degenerescence of metrics towards actions on real trees of Morgan-Shalen [MS84, Ota96].
+The argument follows closely the cases considered by Bonahon-Otal [BO04], Lecuire [Lec06], Anderson-Lecuire [AL13], Baba-Ohshika [BO23] and Lecuire [Lec25].
+The first step is to prove that any sequence of representations with convergent image under $\operatorname{m}^{-}\times\operatorname{b}^{+}$ must have an algebraically convergent subsequence, using the theory of degenerescence of metrics towards actions on real trees of Morgan-Shalen [MS84, Ota96].
 Then, it follows from results of Anderson-Lecuire [AL13] and Bonahon-Otal and Lecuire [BO04, Lec06] that this algebraic limit must lie in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$.
@@
 Combined with the fact that $\operatorname{b}^{+}_{\infty}$ has compact real-analytic fibres and using a theorem of Finney [Fin67], this implies that $\operatorname{b}^{+}_{\infty}$ is injective.
-[…]
-${{[0,\infty)\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{QF}}_{o}(S)}}$${{{\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})}}$${{[0,\infty]\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})}}$${{{\mathcal{AH}}_{o}^{+}(S)}}$${{{\mathcal{ML}}(S^{+})}}$$\scriptstyle{r\times\rm{incl}}$$\scriptstyle{\operatorname{U}}$$\scriptstyle{\operatorname{b}^{+}\times\operatorname{b}^{-}}$$\scriptstyle{\rm{proj}_{2}}$$\scriptstyle{\overline{r}\times\rm{incl}}$$\scriptstyle{\operatorname{B}^{+}}$¯¯$\scriptstyle{\overline{\operatorname{U}}}$$\scriptstyle{\operatorname{b}^{+}}$
+Consider a ray $r\colon[0,\infty)\to{\mathcal{T}}(S^{-})$ such that $\lim_{t\to\infty}r(t)={\mathcal{E}}^{-}$ in the space of end structures ${\mathcal{ES}}(S^{-})$ (called space of gallimaufries in [AL13], see Section 2.8 for a description of its topology), in other words, $r$ extends to a path $\overline{r}\colon[0,\infty]\to{\mathcal{ES}}(S^{-})$ with $\overline{r}(\infty)={\mathcal{E}}^{-}$.
+The objects and maps involved in the proof are summarized in the diagram below, where $\overline{\operatorname{U}}$ denotes the extension of $\operatorname{U}$ to the space of end structures.
+$0,\infty)\times{\mathcal{T}}(S^{+})$${\mathcal{T}}(S^{-})\times{\mathcal{T}}(S^{+})$${\mathcal{QF}}_{o}(S)$${\mathcal{ML}}(S^{-})\times{\mathcal{ML}}(S^{+})$$0,\infty]\times{\mathcal{T}}(S^{+})$${\mathcal{ES}}(S^{-})\times{\mathcal{T}}(S^{+})$${\mathcal{AH}}_{o}^{+}(S)$${\mathcal{ML}}(S^{+})$$\scriptstyle{r\times\rm{incl}}$$\scriptstyle{\operatorname{U}}$$\scriptstyle{\operatorname{b}^{+}\times\operatorname{b}^{-}}$$\scriptstyle{\rm{proj}_{2}}$$\scriptstyle{\overline{r}\times\rm{incl}}$$\scriptstyle{\operatorname{B}^{+}}$$\scriptstyle{\overline{\operatorname{U}}}$$\scriptstyle{\operatorname{b}^{+}}$
 Let $\operatorname{B}^{+}\colon[0,\infty]\times{\mathcal{T}}(S^{+})\to{\mathcal{ML}}(S^{+})$ be the composition $\operatorname{B}^{+}\coloneqq\operatorname{b}^{+}\circ\,\overline{\operatorname{U}}\circ(\overline{r}\times\rm{incl})$, and for each $t\in[0,\infty]$, write $\operatorname{b}_{t}^{+}\coloneqq\operatorname{B}^{+}(t,\cdot)\colon{\mathcal{T}}(S^{+})\to{\mathcal{ML}}(S^{+})$.
@@
 This is expected to hold, and it was proved in the case of rational bending laminations by Bonahon and Otal [BO04].
-But it is not yet known in the general case, where our strategy fails.
+But it is not yet known in the general case, where the current strategy needs to be refined.
 If geometrically finite structures are proven to be parameterized by their bending laminations, then one can prove that the same holds for all structures in ${\mathcal{AH}}_{o}(S)$.
@@
 A transverse measure on a geodesic lamination $L\in{\mathcal{GL}}(S)$ is a map $\mu$ associating to each closed arc transverse to $L$ a number $\mu(k)\in{\mathbb{R}}_{\geq 0}$ (also denoted $i(\mu,k)$) which is invariant under transverse isotopy of arcs, countably additive, and with full support, i.e.
-$\mu(k)>0$ if and only if $k\cap L\not=\emptyset$.
+$\mu(k)>0$ if and only if $k\cap L\not$$\emptyset$.
 Let ${\mathcal{ML}}(S)$ denote the space of measured lamination on $S$, endowed with the weak topology.
@@
 Let $M$ be a hyperbolic three-manifold, $S$ a complete hyperbolic surface (not necessarily compact) and $L$ a geodesic lamination on $S$.
-A pleated surface with pleating locus $L$ is a continuous map $f\colon S\to M$ which is a totally geodesic immersion on $S-L$ and on $L$.
-In general, the flat pieces of $S-L$ are bent along the pleating locus, with an angle recorded into a bending cocycle $\beta$ which is an element of the space ${\mathcal{H}}(L;{\mathbb{R}}/2\pi{\mathbb{Z}})$ of ${\mathbb{R}}/2\pi{\mathbb{Z}}$-valued cocycles along $L$ [Bon96] (which is a torus whose dimension is determined by $S$ and $L$).
+A pleated surface with pleating locus $L$ is a continuous map $f\colon S\to M$ which is a totally geodesic immersion on SLS-L and on $L$.
+In general, the flat pieces of SLS-L are bent along the pleating locus, with an angle recorded into a bending cocycle $\beta$ which is an element of the space ${\mathcal{H}}(L;{\mathbb{R}}/2\pi{\mathbb{Z}})$ of ${\mathbb{R}}/2\pi{\mathbb{Z}}$-valued cocycles along $L$ [Bon96] (which is a torus whose dimension is determined by $S$ and $L$).
 When $f$ is locally convex, this bending cocycle defines a transverse measure with support contained in $L$, called the bending or pleating (measured) lamination of $f$.
@@
 The boundary of the convex core $C(\rho)$ of $\rho\in{\mathcal{AH}}(S)$ is an embedded convex pleated surface [CEM06, II.1.11,12].
-In particular, its bending measured lamination has weight $<\pi$ along closed leaves.
+In particular, its bending measured lamination has weight $\pi$ along closed leaves.
 Now, let $L$ be a maximal geodesic lamination.
@@
 any essential closed curve on $S(e^{-})$ intersects $P\cup L$ or $\operatorname{b}^{+}(\rho)$ (in other words, any essential closed curve on $S$ intersects $P\cup L\cup\Sigma^{-}\cup\operatorname{b}^{+}(\rho)$).
-Hence it lies in the subspace <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ consisting of those measured laminations without closed leaves of weight $\geq\pi$, and which fill $S(e^{-})$ with $e^{-}$.
+Hence it lies in the subspace ${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ consisting of those measured laminations without closed leaves of weight $\geq\pi$, and which fill $S(e^{-})$ with $e^{-}$.
 2.8. Metrics converging to an end structure
-[…]
+In [AL13], Anderson and Lecuire define a natural topology on the space ${\mathcal{ES}}(S)$ of end structures on $S$, i.e.
+of doubly incompressible quadruplets $(P,\Sigma,m,L)$ on $S$ as in the previous section.
+For our purpose we only need to describe one particular type of convergence.
+A sequence of hyperbolic metrics $m_{n}\in{\mathcal{T}}(S)$ (seen as end structures of the form $(\emptyset,\emptyset,m_{n},\emptyset)$) converges to an end structure ${\mathcal{E}}=(P,\Sigma,m,L)$ if the following two conditions hold :
   1. (1)
@@
 there is a non-peripheral simple closed curve $c\subset S(L^{\prime})$ such that $\ell_{m_{n}}(\partial\overline{S(L^{\prime})})/\ell_{m_{n}}(c)\to 0$ and any subsequence of $m_{n}|_{S(L^{\prime})}$ contains a subsequence converging to a projective measured lamination supported by $L^{\prime}$.
-[…]
-[…]
+Thus, we can identify ${\mathcal{T}}(S)$ with an open subspace of ${\mathcal{ES}}(S)$. Note that ${\mathcal{ES}}(S^{-}\sqcup S^{+})={\mathcal{ES}}(S^{-})\times{\mathcal{ES}}(S^{+})$.
+The end invariants of an element $\rho$ of ${\mathcal{AH}}_{o}(S)$ form an end structure $\operatorname{E}(\rho)$ in ${\mathcal{ES}}(S^{-}\sqcup S^{+})$.
+For $\rho\in{\mathcal{AH}}_{o}^{+}(S)$, its end structure $\operatorname{E}(\rho)$ has the form $(P,\Sigma^{-},m^{-},L)\times(\emptyset,\emptyset,m^{+},\emptyset)\in{\mathcal{E}}(S^{-})\times{\mathcal{T}}(S^{+})$ (and any such end structure is doubly incompressible in the sense of [AL13, §2.9]).
 The following is a special case of Theorem D in [AL13].
@@
 This will imply that $\operatorname{b}^{+}_{\infty}\colon{\mathcal{T}}(S^{+})\to{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is proper too.
-The proof is similar to the proof of properness of similar maps by Baba and Ohshika in [BO23] and Anderson and Lecuire in [AL13].
+The proof is very similar to the proof of properness of similar maps by Baba and Ohshika in [BO23], Anderson and Lecuire in [AL13] and Lecuire in [Lec25].
+It essentially follows by combining the last two references.
 Proposition 4.1.
 Let $e^{-}=(P,L)$ be a topological end structure on $S^{-}$. Let $\Sigma^{-}\coloneqq S^{-}-(P\cup S(L))$ be the moderate subsurface of the bottom end, and let $S^{+}=S$ be the one of the top end.
-Let <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ be the subspace of ${\mathcal{ML}}(S^{+})$ consisting of the measured laminations $\mu$ satisfying the following conditions :
+Let ${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ be the subspace of ${\mathcal{ML}}(S^{+})$ consisting of the measured laminations $\mu$ satisfying the following conditions :
   1. (1)
@@
 Proof.
-Write $(\operatorname{m}^{-}\times\operatorname{b}^{+})(\rho_{n})=(m_{n}^{-},\lambda^{+}_{n})$ for each $n\in{\mathbb{N}}$, and $(m^{-},\lambda^{+})$ their limit.
+[…]
 Let us assume that $(\rho_{n})_{n}$ has no algebraically convergent subsequence.
@@
 minimal, maximal and irrational) in its respective component of $S(L)$, hence the support of $\mu$ must be contained in $L$ [Ota96, A.3.10].
-Finally, Bridgeman proved that there is a constant bounding the length of bending laminations of convex cores [Bri98, Prop.
-2] (see §2.4 for the definition of the length of a measured lamination in a Kleinian surface manifold).
-But since $e^{-}$ and $\lambda^{+}$ fill $S(e^{-})$ and $\mu$ has support contained in $L\subset e^{-}$, we obtain that $i(\mu,\lambda^{+})>0$.
-We claim that this forces $(\ell_{\rho_{n}}(\lambda^{+}))_{n}$ to be unbounded, a contradiction with Bridgeman’s result.
-Since $i(\mu,\lambda^{+})>0$, there is a minimal component $\lambda^{+}_{0}\subset\lambda^{+}$ with $i(\mu,\lambda^{+}_{0})>0$.
-If $\lambda^{+}_{0}$ is a closed curve, then (4) implies that $(\ell_{\rho_{n}}(\lambda^{+}_{0}))_{n}$ is unbounded, so is $(\ell_{\rho_{n}}(\lambda^{+}))_{n}$.
-Let us now assume that $\lambda^{+}_{0}$ is irrational.
-We cannot apply (4) to prove the claim since that implication is a priori only valid for closed curves (and, by linearity, for weighted multicurves).
-Better control on (4) is given by Otal’s continuity theorem [Ota96, Theorem 4.0.1], of which we need the more general form stated by Lecuire in [Lec20, Theorem 3.7].
+Since $e^{-}$ and $\lambda^{+}$ fill $S(e^{-})$ and $\mu$ has support contained in $L\subset e^{-}$, we obtain that $i(\mu,\lambda^{+})>0$.
+Thus there is a minimal component $\lambda^{+}_{0}\subset\lambda^{+}$ with $i(\mu,\lambda^{+}_{0})>0$.
+In other words, the support of $\lambda^{+}_{0}$ is realized in ${\mathcal{T}}_{\mu}$.
+This leads to a contradiction, as Lecuire shows in [Lec25].
+His argument is divided into two cases, depending on whether $\lambda^{+}_{0}$ is supported on an irrational lamination [Lec25, §3.1] or a closed curve [Lec25, §3.2].
+The resulting contradictions stem from Otal’s continuity theorem applied to $\lambda^{+}_{0}$ [Ota96, Theorem 4.0.1][Lec20, Theorem 3.7], which provides uniform control on the growth of lengths of curves (or weighted multicurves, by linearity) that are close to a lamination realized in the limit tree.
+Using convergence of the bending laminations $(\lambda^{+}_{n})_{n}$ to $\lambda^{+}$, Lecuire constructs sequences of weighted simple closed curves whose lengths with respect to $\rho_{n}$ grow too slowly, leading to a contradiction.
 Theorem 4.4 (Otal continuity theorem).
@@
 each leaf of $\alpha$ intersects $\mu$ transversely.
-Then there exists a neighborhood ${\mathcal{V}}(\alpha)\subset{\mathcal{GL}}(S)$ of $\alpha$ and constants $K$, $n_{0}$ such that, for any simple closed curve $c\subset{\mathcal{V}}(\mu)$ and for any $n\geq n_{0}$,
+Then there exists a neighborhood ${\mathcal{V}}(\alpha)\subset{\mathcal{GL}}(S)$ of $\alpha$ and constants $K>0$, $n_{0}$ such that, for any simple closed curve $c\in{\mathcal{V}}(\alpha)$ and for any $n\geq n_{0}$,
 (5) $\epsilon_{n}\ell_{\rho_{n}}(c)\geq K\ell_{s_{0}}(c).$
-Pick a sequence of weighted simple closed curves $(\alpha_{j})_{j\in{\mathbb{N}}}$ converging to $\lambda^{+}_{0}$ in ${\mathcal{ML}}(S)$.
-By Otal’s continuity theorem there exist constants $K$, $N$ and $J$ such that, for all $n\geq N$ and $j\geq J$,
-$\ell_{\rho_{n}}(\alpha_{j})\geq\frac{K}{\epsilon_{n}}\ell_{s_{0}}(\alpha_{j}),$
-where $s_{0}$ is any fixed reference metric, independent of $n$ and $j$. By continuity of length functions (Theorem 2.6 by Brock [Bro00]), the limit $j\to\infty$ gives, for each $n\geq N$,
-$\ell_{\rho_{n}}(\lambda^{+}_{0})\geq\frac{K}{\epsilon_{n}}\ell_{s_{0}}(\lambda^{+}_{0}).$
-Since $\epsilon_{n}\searrow 0$, while $K\ell_{s_{0}}(\lambda^{+}_{0})$ is a positive constant, we obtain that $\ell_{\rho_{n}}(\lambda^{+}_{0})\to\infty$, concluding the proof. ∎
+[…]
+In the second case, $\lambda_{0}^{+}$ is supported on a simple closed curve $d$.
+In particular, $\epsilon_{n}\ell_{\rho_{n}}(d)\to i(\mu,d)>0$.
+In that case, Lecuire shows that any closed curve $c$ transverse to $d$ satisfies [Lec25, Lemma 3.5]
+$\frac{\epsilon_{n}\ell_{\rho_{n}}(d)}{\epsilon_{n}\ell_{\rho_{n}}(c)}=\frac{\ell_{\rho_{n}}(d)}{\ell_{\rho_{n}}(c)}\to 0.$
+But $\epsilon_{n}\ell_{\rho_{n}}(c)\to i(\mu,c)$, so $\epsilon_{n}\ell_{\rho_{n}}(d)\to 0$, a contradiction. This concludes the proof. ∎
 4.5. Geometric convergence
-Let $(\rho_{n})_{n}$ be a sequence of representations in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ as in Proposition 4.3 with $\operatorname{m}^{-}\times\operatorname{b}^{+}(\rho_{n})=(m^{-}_{n},\lambda^{+}_{n})$ converging to $(m^{-},\lambda^{+})$ in ${\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$.
-By Proposition 4.3, up to replacing $(\rho_{n})_{n}$ by a subsequence, we assume that it converges algebraically to a representation $\rho_{\infty}\in{\mathcal{AH}}_{o}(S)$.
-The goal of this section is to show that $\rho_{\infty}$ belongs to ${\mathcal{AH}}_{o}^{+}(S;e^{-})$, i.e.
-that ${\mathbb{H}}^{3}/\rho_{\infty}(\Gamma)$ has parabolic locus $P$, no new geometrically infinite end, and that it still has $L$ as its ending lamination.
+[…]
 First, the bottom end is controlled.
@@
 The moderate part of the statement is [AL13, Lemma 4.1] and the immoderate part is [AL13, Lemma 4.2]. ∎
-The top end can be controlled thanks to the following result of Bonahon-Otal [BO04, §5,6] and Lecuire [Lec04, §4.1], [Lec06, §4.2].
+The top end can be controlled thanks to the following result of Bonahon-Otal [BO04, §5,6] and Lecuire [Lec25, §4.1], [Lec06, §4.2].
 We state a slightly weaker version, although their result also works in presence of parabolics in the top end.
@@
 Let $(\lambda_{n})_{n}$ be the sequence of bending laminations of those pleated surfaces.
-Assume that $\lambda_{n}\to\lambda$ in ${\mathcal{ML}}(S)$ and that the closed leaves of $\lambda_{n}$ and $\lambda$ have weight $<\pi$.
+Assume that $\lambda_{n}\to\lambda$ in ${\mathcal{ML}}(S)$ and that the closed leaves of $\lambda_{n}$ and $\lambda$ have weight $\pi$.
 Then $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ converges geometrically to a convex pleated surface $(f_{\infty},\pi_{1}(S),\rho_{\infty})$ inside ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$.
@@
 By Bonahon’s shear-bend parameterization Theorem 2.5, the open subset ${\mathcal{R}}(L)\subset{\mathcal{X}}(S)$ consisting of those representations realizing a pleated surface with pleating locus contained in $\lambda^{+}$ is biholomorphically homeomorphic to an open subset of ${\mathcal{H}}(L;{\mathbb{C}}/2\pi i{\mathbb{Z}})$ via the shear-bend coordinates $\sigma+i\beta$.
-For a generic $\rho\in{\mathcal{R}}(L)$, the associated pleated surface $P_{\rho}$ lies inside the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ and is neither convex nor concave, in particular $\beta(\rho)$ takes positive and negative values in $[-\pi,\pi]$, i.e.
+For a generic $\rho\in{\mathcal{R}}(L)$, the associated pleated surface $P_{\rho}$ lies inside the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ and is neither convex nor concave, in particular $\beta(\rho)$ takes positive and negative values in $\pi,\pi]$, i.e.
 it does not correspond to a transverse measure.
@@
 6.2. Prescribed third fundamental form
-Given $\rho\in{\mathcal{QF}}_{o}(S)$, a theorem of Labourie [Lab92] states that each end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ is uniquely foliated by smooth closed subsurfaces of constant curvature equal to $K$, with $K$ contained in the interval $(-1,0)$.
+Given $\rho\in{\mathcal{QF}}_{o}(S)$, a theorem of Labourie [Lab92] states that each end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ is uniquely foliated by smooth closed subsurfaces of constant curvature equal to $K$, with $K$ contained in the interval (1,0)(-1,0).
 The third fundamental form of such a subsurface has constant curvature $K^{*}\coloneqq K/(K+1)$.
@@
 Focusing on the top end, it follows that, for each $K\in(-1,0)$, one can define the map
-(6) $III^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}^{K^{*}}(S)$
+(6) $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}^{K^{*}}(S)$
 sending a quasifuchsian representation $\rho$ to the third fundamental form of the constant curvature $K$ surface in the top end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$.
-The third fundamental form $III^{+}_{K}$ is a “smooth measure of bending” and Chen-Schlenker proved that the analogous “smooth version” of Proposition 6.1 holds (see Corollary 1.4 and the proof of Lemma 4.1 in [CS25]).
+The third fundamental form $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$ is a “smooth measure of bending” and Chen-Schlenker proved that the analogous “smooth version” of Proposition 6.1 holds (see Corollary 1.4 and the proof of Lemma 4.1 in [CS25]).
 We need the following special case of their result, which is more general.
 Theorem 6.3 (Chen-Schlenker [CS25]).
-The map $\operatorname{m}^{-}\times III^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{T}}^{K^{*}}(S^{+})$ is a homeomorphism.
-This is a mixed version of the earlier result of Schlenker that $III^{-}_{K}\times III^{+}_{K}$ is a homeomorphism [Sch06], which was an essential ingredient in [DS24].
+The map $\operatorname{m}^{-}\times I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{T}}^{K^{*}}(S^{+})$ is a homeomorphism.
+This is a mixed version of the earlier result of Schlenker that $I\hskip-2.84544ptI\hskip-2.84544ptI^{-}_{K}\times I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$ is a homeomorphism [Sch06], which was an essential ingredient in [DS24].
 These smooth analogues of $\operatorname{m}^{-}\times\operatorname{b}^{+}$ actually approximate $\operatorname{m}^{-}\times\operatorname{b}^{+}$, as follows from the following result of Bonsante-Mondello-Schlenker [BMS13] and Belraouti [Bel17].
@@
 Proposition 6.4.
-When $K\to-1$, the maps $III^{+}_{K}$ converge to $\operatorname{b}^{+}$ in the length spectrum, uniformly on compacts.
-More precisely, $\ell_{III^{+}_{K}(\rho)}(c)\to i(\lambda^{+}(\rho),c)$ when $K\to-1$, for each closed curve $c$ in $S$ and each $\rho\in{\mathcal{QF}}_{o}(S)$, and the convergence is uniform for $\rho$ in a compact subset of ${\mathcal{QF}}_{o}(S)$.
-In order to apply the variant of Finney’s theorem as in the next section, we identify the targets of the maps $III^{+}_{K}$, for $K\in(-1,0)$, and $\operatorname{b}^{+}$, via the earthquake map $E\colon{\mathcal{ML}}(S)\times{\mathcal{T}}(S)\to{\mathcal{T}}(S)\colon(\lambda,m)\mapsto E(\lambda,m)$ as in [DS24, §4.1].
+When $K\to-1$, the maps $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$ converge to $\operatorname{b}^{+}$ in the length spectrum, uniformly on compacts.
+More precisely, $\ell_{I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}(\rho)}(c)\to i(\lambda^{+}(\rho),c)$ when $K\to-1$, for each closed curve $c$ in $S$ and each $\rho\in{\mathcal{QF}}_{o}(S)$, and the convergence is uniform for $\rho$ in a compact subset of ${\mathcal{QF}}_{o}(S)$.
+In order to apply the variant of Finney’s theorem as in the next section, we identify the targets of the maps $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$, for $K\in(-1,0)$, and $\operatorname{b}^{+}$, via the earthquake map $E\colon{\mathcal{ML}}(S)\times{\mathcal{T}}(S)\to{\mathcal{T}}(S)\colon(\lambda,m)\mapsto E(\lambda,m)$ as in [DS24, §4.1].
 More precisely, we fix a point $m_{0}\in{\mathcal{T}}(S)$ and consider, for each $K\in(-1,0)$, the map
-(7) $V_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{+})\colon\rho\mapsto\frac{1}{\sqrt{\left|K^{*}\right|}}E(-,m_{0})^{-1}(\left|K^{*}\right|III^{+}_{K}(\rho)),$
+(7) $V_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{+})\colon\rho\mapsto\frac{1}{\sqrt{\left|K^{*}\right|}}E(-,m_{0})^{-1}(\left|K^{*}\right|I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}(\rho)),$
 which is such that $\operatorname{m}^{-}\times V_{K}$ is a homeomorphism since the earthquake map is a homeomorphism by Thurston’s earthquake theorem [Thu86].
@@
 6.6. Fibres of a limit of homeomorphism
-Let $\operatorname{\mathbb{S}}^{k}\subset{\mathbb{D}}^{k+1}$ denote the inclusion of the $k$-sphere as the boundary of the $(k+1)$-ball.
-Recall that a space $K$ is weakly contractible if, for any $k\geq 0$, any map $\operatorname{\mathbb{S}}^{k}\to K$ extends to a map ${\mathbb{D}}^{k+1}\to K$.
-If a space $X$ is weakly contractible, then it is acyclic, i.e.
-all its reduced homology groups are trivial.
 A subset $K\subset N$ is a neighbourhood retract if it has an open neighbourhood $\Omega\subset N$ such that the inclusion $j\colon K\xhookrightarrow{}\Omega$ has a retraction $r\colon\Omega\to K$, i.e.
@@
 $f_{n}(x_{n})\to f_{\infty}(x_{\infty})$ whenever $x_{n}\to x_{\infty}$.
-If a fibre of $f_{\infty}$ is compact and is an absolute neighbourhood retract, then it is weakly contractible.
-In fact, a compact weakly contractible ANR is automatically contractible, although we only need the fact that it is acyclic.
+If a fibre of $f_{\infty}$ is compact and is an absolute neighbourhood retract, then it is contractible.
 Proof.
-Let $K\coloneqq f_{\infty}^{-1}(\ast)$ be a compact fibre, which is a retract of an open neighbourhood $\Omega$, which we can assume to have compact closure.
-Assuming that $K$ is not weakly contractible, we will show that $f_{\infty}$ sends a point of $\partial\Omega$ to $\ast$.
-By contradiction, assume that there exists $k\geq 0$ and a map $\sigma\colon S^{k}\to K$, which is not contractible, i.e.
-does not extend to a map ${\mathbb{D}}^{k+1}\to K$.
-In particular, there is also no map ${\mathbb{D}}^{k+1}\to\Omega$ extending $\sigma$, for one could compose it with a retraction $\Omega\to K$ to obtain a map which lands in $K$.
+Let $K\coloneqq f_{\infty}^{-1}(\ast)$ be a (nonempty) compact fibre, which is a retract of an open neighbourhood $\Omega$, which we can assume to have compact closure.
+Let $j\colon K\xhookrightarrow{}\Omega$ be the inclusion and $r\colon\Omega\to K$ a retraction.
+Assuming that $K$ is not contractible, we will show that $f_{\infty}$ sends a point of $\partial\Omega$ to $\ast$.
 Let $U\subset N^{\prime}$ be a contractible neighbourhood of $\ast$.
@@
 Fix such a $n\in{\mathbb{N}}$.
-Thus, $f_{n}(K)\subset U$.
-Since $U$ is contractible, the map $f_{n}\circ\sigma\colon\operatorname{\mathbb{S}}^{k}\to U$ extends to a map $g\colon{\mathbb{D}}^{k+1}\to U$.
-But recall that $f_{n}$ is an open embedding, hence $f_{n}\circ\sigma$ is not contractible inside $f_{n}(\Omega)$.
-Thus $g({\mathbb{D}}^{k+1})$ cannot be entirely contained in $f_{n}(\Omega)$, so it must intersect $f_{n}(\partial\Omega)$, say at a point $f_{n}(x_{U})$, for otherwise $g({\mathbb{D}}^{k+1})$ would be disconnected.
-In particular, $x_{U}\in\partial\Omega$ and $f_{n}(x_{U})\in U$.
+Thus $f_{n}(K)\subset U$.
+Since $U$ is contractible, there is a homotopy $H\colon K\times[0,1]\to U$ from $H(-,0)=f_{n}|_{K}\colon K\to U$ to a constant map H(,1)H(-,1).
+But recall that $f_{n}$ is an open embedding.
+Thus $H(K\times[0,1])$ cannot be entirely contained in $f_{n}(\Omega)$, for otherwise $r\circ(f_{n}|_{\Omega})^{-1}\circ H\colon K\times[0,1]\to K$ would be a nulhomotopy of $\operatorname{id}_{K}$.
+Since $H(K\times[0,1])$ is connected ($H$ is a nulhomotopy, hence every point in its image is connected by a path to the constant image of H(,1)H(-,1)), it must intersect $f_{n}(\partial\Omega)$ (which equals $\partial f_{n}(\Omega)$ since $f_{n}$ is an open embedding), say at a point $f_{n}(x_{U})$ with $x_{U}\in\partial\Omega$.
+In particular, $f_{n}(x_{U})\in U$.
 Considering a nested sequence of contractible neighbourhoods $U_{0}\supset U_{1}\supset\dots\ni\ast$ such that $\bigcap_{i=0}^{\infty}U_{i}=\ast$, the above paragraph gives a sequence $(n_{i})_{i\in{\mathbb{N}}}$ in $\mathbb{N}$}, with $n_{i}\to\infty$, and a sequence of points $(x_{i})_{i\in{\mathbb{N}}}$ in $\partial\Omega$, such that $f_{n_{i}}(x_{i})\in U_{i}$ for all $i\in{\mathbb{N}}$.
@@
 $f_{\infty}(x_{\infty})=\lim_{i\to\infty}f_{n_{i}}(x_{i})=\ast,$
-which is impossible since $\partial\Omega$ does not meet $K=f_{\infty}^{-1}(\ast)$. This concludes the proof that $K$ is weakly contractible. ∎
+which is impossible since $\partial\Omega$ does not meet $K=f_{\infty}^{-1}(\ast)$. This concludes the proof that $K$ is contractible. ∎
 We are finally able to prove Proposition 6.1, i.e.
@@
 By Lemmas 5.1 and 5.2, $F$ is a real-analytic space and an ANR.
-Since $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is the limit of the homeomorphisms $\operatorname{m}^{-}\times V_{K}$ and the convergence is uniform on compacts, $F$ is weakly contractible by Theorem 6.7, hence it is acyclic.
+Since $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is the limit of the homeomorphisms $\operatorname{m}^{-}\times V_{K}$ and the convergence is uniform on compacts, $F$ is contractible by Theorem 6.7.
 Since $F$ is a compact real-analytic space, it has a fundamental class modulo two (by Borel-Haefliger and Sullivan, see Theorem 5.3), i.e.
 a nonzero homology class $\alpha\in H_{\dim(F)}(F;{\mathbb{Z}}/2)$.
-But $F$ is acyclic (in fact, contractible), hence its homology is trivial in all degrees except 00.
+But $F$ is contractible, hence its homology is trivial in all degrees except 0.
 Since $\alpha\neq 0$, we must have $\dim(F)=0$.
-Therefore, $F$ is a contractible compact real-analytic space of dimension 00, that is, a point.
+Therefore, $F$ is a contractible compact real-analytic space of dimension 0, that is, a point.
 This proves that $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is injective.
@@
 $\operatorname{m}^{-}\times\operatorname{b}^{+}\colon{\mathcal{AH}}_{o}^{+}(S;e^{-})\to{\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$
-is proper (Proposition 4.1), continuous (on the first factor, by the simultaneous uniformization theorem, see Section 2.10 ; on the second factor, by Proposition 3.4), and its domain and target are manifolds of the same dimension (By [Ker92, Theorem 4.7], <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is a manifold of the same dimension as ${\mathcal{T}}(S^{+})$), it suffices to show that it is injective.
+is proper (Proposition 4.1), continuous (on the first factor, by the simultaneous uniformization theorem, see Section 2.10 ; on the second factor, by Proposition 3.4), and its domain and target are manifolds of the same dimension (By [Ker92, Theorem 4.7], ${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is a manifold of the same dimension as ${\mathcal{T}}(S^{+})$), it suffices to show that it is injective.
 With the notation of Section 1.1, it suffices to show that $\operatorname{b}^{+}_{\infty}$ is injective.